### v5

# Conditional Denoising Meets Polynomial Modeling: A Flexible Decoupled Framework for Time Series Forecasting

Abstract
Time series forecasting models are becoming increasingly prevalent due to their critical role in decision-making across various domains. However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between its specific components. In particular, fluctuating patterns and smooth trends within time series exhibit distinct characteristics. In this work, to model complicated temporal patterns, we propose a Conditional Denoising Polynomial Modeling (CDPM) framework, where probabilistic diffusion models and deterministic linear models are trained end-to-end. Instead of modeling the coupled time series, CDPM decomposes it into trend and seasonal components for modeling them separately. To capture the fluctuating seasonal component, we employ a probabilistic diffusion model based on statistical properties from the historical window. For the smooth trend component, a module is proposed to enhance linear models by incorporating historical dependencies, thereby preserving underlying trends and mitigating noise distortion. Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models. [^2] Our code is open-sourced and available at https://github.com/zjt-gpu/FDF.
1 Introduction
Time series forecasting is a fundamental task across various domains, playing a crucial role in driving informed decision-making in industries such as finance, healthcare, and manufacturing Kashpruk et al. (2023); Liu et al. (2024). It involves analyzing historical patterns to model the dynamic evolution of underlying processes and predict future trends Montgomery et al. (2015). Accurate time series forecasting is essential for tasks such as demand prediction, resource allocation, enabling organizations to anticipate future needs and optimize operations. Trastour et al. (2002); Kilcioglu et al. (2017); Drutsa et al. (2015)
Figure 1: The forward diffusion process iteratively adds noise to the original time series, which causes the trend components to be overshadowed as the noise increasingly dominates. This makes it challenging to recover trend during the reverse denoising process.
Over the years, various approaches have been proposed to address this challenge, ranging from classical statistical models, such as Autoregressive Integrated Moving Average (ARIMA) Asteriou and Hall (2011), to machine learning techniques Masini et al. (2023), and advanced deep learning models, including recurrent neural networks (RNNs) Lin et al. (2023); Lai et al. (2018), convolutional neural networks (CNNs) Bai et al. (2018); Koprinska et al. (2018), and transformer-based models Zhou et al. (2021); Wu et al. (2021). Large Language Models (LLM) like Time-LLM and LLMTIME Jin et al. (2023); Gruver et al. (2024) have demonstrated the potential in time series forecasting. In addition to their success, probabilistic models have emerged as promising alternatives, offering a principled framework for quantifying uncertainty and modeling complex distributions. Diffusion models Rasul et al. (2021); Shen and Kwok (2023), becoming the leading paradigm for probabilistic models originally developed for generation, have proven effective in time series forecasting due to their ability to capture intricate stochastic processes.
Diffusion-based methods have shown significant potential in time series forecasting. We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fail to account for the distinct dynamics of different patterns. Specifically, noise is uniformly added across the entire time series, and denoising is then applied indiscriminately, treating all temporal components equally. However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require distinct treatment. Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties. A uniform noise addition and denoising strategy fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting. As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing forecasting accuracy.
Employing suitable models to capture the different components is crucial. Single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s forecasting ability. Diffusion models, with their advantages in modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series. In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends. Therefore, by leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more comprehensively capture the dynamic variations within the series and enhance forecasting accuracy.
To overcome these challenges, we propose a novel Conditional Denoising Polynomial Modeling (CDPM) framework that leverages the inductive bias of time series decomposition. By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to each component’s characteristics. Specifically, we employ carefully designed Conditional Denoising Seasonal Module (CDSM) to capture the high-frequency seasonal component, introducing an innovative denoising strategy conditioned on statistical properties derived from historical window. Concurrently, we model the intrinsic trend component using Polynomial Trend Module (PTM), which are adept at capturing gradual changes over time. Additionally, we reformulate the Improved Evidence Lower Bound (IELBO) to facilitate the joint training of the CDSM and PTM. This decoupling method enhances the framework’s ability to capture temporal dependencies and complex patterns. The key contributions of this work are:
  • We introduce a novel framework that enables the decoupled modeling of trend and seasonal components, providing new insights into time series forecasting.
  • We design tailored methods and reformulate ELBO for each component-using CDSM for the seasonal component and PTM for the trend component—to effectively capture their distinct behaviors.
  • We conduct comprehensive experiments on benchmark datasets to validate effectiveness of our framework, we also present the rationale for modeling different components with several case studies.
Figure 2: Illustration of the proposed Conditional Denoising Polynomial Modeling (CDPM) framework. The Conditional Denoising Seasonal Module (CDSM) refines the seasonal component through a diffusion process conditioned on historical patterns, while the Polynomial Trend Module (PTM) models the trend component to capture different patterns.
2 RELATED WORK
2.1 Deterministic Forecasting
Deterministic models map deterministic inputs to deterministic outputs by learning consistent patterns within time series, making them crucial for capturing predictable temporal relationships. Recurrent neural networks (RNNs), such as LSTNet Lai et al. (2018) and SegRNN Lin et al. (2023), became popular for sequential dependencies, while convolutional neural networks (CNNs) like TCN Bai et al. (2018) and SCINet Liu et al. (2022) excelled at capturing both local and global temporal features. Graph-based models, such as MTGNN Wu et al. (2020) and FourierGNN Yi et al. (2024), improved forecasting by modeling interdependencies between time series. More recently, Transformer-based models like Informer Zhou et al. (2021) and Autoformer Wu et al. (2021) have utilized self-attention mechanisms to efficiently capture long-range dependencies, excelling in high-dimensional and long-horizon tasks. Surprisingly, simpler models like DLinear Zeng et al. (2023) and TSMixer Ekambaram et al. (2023) have shown that well-tuned linear layers, focusing on long-term trends, can outperform complex frameworks, highlighting their efficiency. Despite advances, deterministic models often struggle with volatile fluctuations, limiting their ability to model complex time series dynamics.
2.2 Probabilistic Forecasting
Probabilistic models offer uncertainty estimation and the ability to model temporal distributions. These models, including Variational Autoencoders (VAEs)Kingma and Welling (2022), Generative Adversarial Networks (GANs)Goodfellow et al. (2014), and flow-based models Rezende and Mohamed (2016), are designed to capture high-dimensional distribution. Recently, Denoising Diffusion Probabilistic Models (DDPMs) have emerged as a promising paradigm for generative modeling, demonstrating their potential in various time series applications. DDPMs have advanced time series forecasting by capturing complex temporal patterns. Early models like TimeGrad Rasul et al. (2021) introduced autoregressive denoising with Langevin sampling for multivariate predictions. TSDiff Kollovieh et al. (2024) improved short-term accuracy through self-guiding mechanisms, while score-based models such as ScoreGrad Yan et al. (2021) applied stochastic differential equations (SDEs) for continuous-time forecasting, expanding diffusion to fluctuatingly sampled data. Conditional models like TimeDiff Shen and Kwok (2023) incorporated external information to guide the diffusion process, outperforming traditional methods. Latent Diffusion Models (LDMs)Rombach et al. (2022) enhanced efficiency by performing diffusion in lower-dimensional spaces. This approach was further demonstrated by LDCast and Latent Diffusion Transformers (LDT)Feng et al. (2024), which improved precipitation and time series forecasting. Models like DSPD and CSPD Bilos et al. (2023) applied diffusion to anomaly detection and interpolation. Diffusion models have also been successfully used in flood forecasting Shao et al. (2024), stock market prediction Daiya et al. (2024), and EV load forecasting Li et al. (2024).
Despite the advancements in both deterministic and probabilistic forecasting approaches, existing research typically applies each paradigm separately to represent time series, with limited efforts to integrate them into a unified modeling framework. However, effectively capturing the diverse characteristics of time series, such as the smooth trend and fluctuating seasonal components, requires leveraging both deterministic and probabilistic models.
3 Notations and Problem Statement
Time series represent numerical sequence ordered chronologically, typically recorded at regular temporal intervals. Formally, consider a multivariate time series $X_{L+T}\in\mathbb{R}^{(L+T)\times d}$, where $L$ and $T$ denotes the length of the historical and the length of the predicted window respectively, with $d$ channels recorded for each time step. The objective of time series forecasting is to train a model $G$ that learns to map the historical series $X_{L}$ to the predicted series $X_{T}$. The model leverages the information embedded in the historical series to predict future series, aiming to capture both temporal dependencies for accurate forecasting.
4 METHODOLOGY
In this section, we formalize the time series forecasting problem and present the proposed Conditional Denoising Polynomial Modeling (CDPM) framework, comprising two core components: the Conditional Denoising Seasonal Module (CDSM) and the Polynomial Trend Module (PTM).
4.1 Overall Framework
The CDPM addresses time series forecasting by independently modeling trend and seasonal patterns, thereby capturing distinct temporal behaviors more effectively. As depicted in Fig. 2, CDPM consists of two main modules: CDSM and PTM. The CDSM employs a denoising diffusion process conditioned on statistical properties from the historical window to accurately capture and refine seasonal variation patterns. In parallel, the PTM models trends using polynomial linear layers, providing flexibility in handling both linear and nonlinear trends. By decoupling and specifically modeling these components, CDPM enhances forecasting accuracy, offering an effective solution for capturing the complex temporal dynamics inherent in time series. We also reformulate the Evidence Lower Bound (ELBO) to better jointly train the PTM and CDSM. By handling trend and seasonal components independently, CDPM allows each module to focus on its respective dynamics, leading to enhanced overall performance.
4.2 Preprocessing for CDPM
Instance Normalization.
Instance normalization across both historical and predicted windows is essential for maintaining the stability of the CDPM. We normalize the historical window by subtracting the mean and dividing by the standard deviation Kim et al. (2021), and transfer the same statistics to normalize the predicted window during training. Specifically, $x_{L}$ is normalized to $x_{L}^{0}$, and $x_{T}$ is normalized to $x_{T}^{0}$ using the statistics derived from $x_{L}$. This ensures alignment between the windows and preserves both the temporal structure and stability in the forecasting process. The subsequent denormalization step restores the forecasted values to their original distribution, thereby maintaining consistency with input data.
Decomposition Operation.
Decomposition is an important approach to achieving decoupling in time series analysis, as it allows for the separation of trend components and seasonal components Cleveland et al. (1990), enabling each characteristic to be handled independently. We decompose both historical $X_{L}^{0}$ and predicted $X_{T}^{0}$ windows into trend and seasonal components. The trend component $X_{L,t}^{0}$ of historical window is estimated using a moving average:
$X_{L,t}^{0}=\text{Avg}(X_{L}^{0}),$ (1)
where Avg denotes the moving average.
$X_{L,s}^{0}=X_{L}^{0}-X_{L,t}^{0},$ (2)
where the seasonal component $X_{L,s}^{0}$ is the residual after subtracting the trend. Thus, we obtain the trend component $X_{L,t}^{0}$ and seasonal component $X_{L,s}^{0}$ of historical window.
Similarly, we compute $X_{T,t}^{0}$ and $X_{T,s}^{0}$ for predicted window. Decomposition operation isolates the distinct characteristics of time series.
4.3 Conditional Denoising Seasonal Module
Modeling fluctuating seasonal components in time series is challenging due to their non-stationary and intricate patterns. We propose the CDSM, based on denoising diffusion model, which is a popular paradigm in probabilistic modeling. By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
Forward Diffusion Process.
In the forward diffusion process, noise is progressively added to the seasonal component. At each diffusion step $k$, the noisy seasonal series is:
$X_{T,s}^{k}=\sqrt{\bar{\alpha}_{k}}\,x_{T,s}^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon,$ (3)
where $X_{T,s}^{k}$ is the seasonal component at step $k$, $x_{T,s}^{0}$ is the original seasonal component, $\epsilon\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\bar{\alpha}_{k}\in[0,1]$ is the noise schedule. As $k$ increases, the seasonal component becomes more noise-dominated.
Reversed Conditional Denoise.
The denoising process starts by embedding $X_{T,s}^{k}$:
$\mathbf{h}=\text{Emb}(X_{T,s}^{k}),$ (4)
where $\text{Emb}(\cdot)$ is a convolutional MLP that maps the noisy sequence to a hidden dimension while capturing local temporal features. Sinusoidal positional encodings for diffusion step $t$ are integrated via Adaptive Layer Normalization (AdaLN):
$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t}.$ (5)
where $\alpha_{t}$ and $\beta_{t}$ are learnable parameters projected from the diffusion step encoding.
The decoder composed of convolutional layers and MLPs, captures global dependencies and feature interactions. Convolutional layers extract fine-grained local temporal variations, while MLPs model cross-dimensional relationships. Layer normalization and a projection layer map the output back to the original feature space.
To enhance denoising process, we incorporate conditional information from the historical seasonal component $X_{L,s}^{0}$.
$\mu_{L,i}=\frac{1}{H}\sum_{h=1}^{H}x_{L,s,i,h},$ (6)
$\sigma_{L,i}^{2}=\frac{1}{H}\sum_{h=1}^{H}\left(x_{L,s,i,h}-\mu_{L,i}\right)^{2},$ (7)
where, for each patch $P_{L,i}$, $\mu_{L,i}$ represents the local mean and $\sigma_{L,i}^{2}$ denotes the variance. The historical data is divided into patches $P$ of length $H$, with each patch capturing localized seasonal signals.
These statistics capture seasonal variability. The predicted mean $\hat{\mu}_{T,i}$ and variance $\hat{\sigma}_{T,i}^{2}$ for the target window $P_{T,i}$ are computed using two MLPs:
$\hat{\mu}_{T,i}=\text{MLP}_{\mu}(\mu_{L,i}),$ (8)
$\hat{\sigma}_{T,i}^{2}=\text{MLP}_{\sigma}(\sigma_{L,i}^{2}),$ (9)
where $\text{MLP}_{\mu}(\cdot)$ and $\text{MLP}_{\sigma}(\cdot)$ capture the mean and variance of historical window features, respectively. The predicted statistics are used in the denoising process via Gaussian sampling:
$\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{T,i}+\hat{\sigma}_{T,i}\cdot z,$ (10)
where $z\sim\mathcal{N}(0,I)$. This enables the model to incorporate stochastic variability while aligning with historical patterns.
Finally, the denoised output $\hat{x}_{T,s}^{0}$ is combined with the conditional information through a weighted summation:
$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1}.$ (11)
where $\rho_{1}$ and $\rho_{2}$ are learnable parameters that balance the denoised series and conditional information, ensuring effective seasonal modeling while maintaining statistical properties.
4.4 Polynomial Trend Module
Different trends exhibit distinct patterns of progression, including periods of growth and decline, making it essential to capture a broader range of trend behaviors. By leveraging Cover’s theorem Cover (1965), we project smooth time series patterns into a polynomial space, facilitating the extraction of both linear and nonlinear trends. Based on this, polynomial modeling Liu et al. (2021) is employed to better capture these diverse trend variations. To fully harness polynomial modeling, we introduce a square root transformation as a nonlinear transformation within the polynomial framework. The square root transformation effectively smooths extreme values, enhancing the framework’s flexibility in capturing a wide range of trend patterns.
As a deterministic model, PTM consists of two complementary pathways. The first models stable trends with a linear layer applied directly to the historical trend series:
$T_{\text{origin}}(x_{T,t}^{0})=\text{Linear}_{\text{origin}}(x_{T,t}^{0}),$ (12)
where $\text{Linear}_{\text{origin}}(\cdot)$ captures the general trend.
The second pathway enhances trend modeling by applying square root transformation, thereby mitigating the impact of extreme values and stabilizing trend predictions:
$T_{\text{sqrt}}(x_{T,t}^{0})=\text{Linear}_{\text{root}}(\sqrt{x_{T,t}^{0}}),$ (13)
where $\text{Linear}_{\text{sqrt}}(\cdot)$ captures the transformed trend.
The outputs from both pathways are combined via weighted summation to capture a wider range of trends:
$f_{\theta}(x_{T,t}^{0})=\lambda_{1}T_{\text{origin}}(x_{T,t}^{0})+\lambda_{2}T_{\text{sqrt}}(x_{T,t}^{0}),$ (14)
where $\lambda_{1}$ and $\lambda_{2}$ are learnable parameters that balance the contributions of the linear and root-based components. This mechanism allows the model to adapt to both linear and sub-linear growth patterns, enhancing its ability to forecast complex time series with diverse dynamics.
4.5 Reconstruction and Optimization
Integrating trend and seasonal components improves forecasting accuracy by combining their strengths for more comprehensive predictions. Finally, the derived loss function enables joint optimization, allowing both components to be learned concurrently and effectively.
Reconstruction.
During the training and inference phases, after the trend component $\hat{x}_{T,t}^{0}$ is predicted by the PTM and the seasonal component $\hat{x}_{T,s}^{0}$ is recovered by the CDSM.
The final prediction $\hat{x}_{T}^{0}$ calculated by integrating the trend and seasonal components as follows:
$\hat{x}_{T}^{0}=X_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{s,t}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\mid\nabla_{\mathbf{x}_{T,s}}^{k}\right)+f_{\theta}(x_{T,t}^{0}),$ (15)
where $X_{\theta}(\cdot)$ represents the CDSM and $f_{\theta}(\cdot)$ corresponds to the PTM.
Improved Evidence Lower Bound Optimization.
The final prediction $\hat{x}_{T}^{0}$ is subsequently denormalized to $\hat{x}_{T}$ using the mean and variance derived from the historical window, ensuring that the forecast is consistent with the original distribution of the time series. The loss function in CDPM is derived from the Evidence Lower Bound (ELBO). Furthermore, we extend the traditional ELBO to accommodate the separate modeling of the trend and seasonal components. This decoupled modeling of the trend and seasonal components ensures that both contribute independently to the final forecast. Through derivation, we find that by modeling the seasonal component with CDSM and the trend component with PTM, the results can be jointly optimized and trained to improve the overall performance of our framework, leading to more accurate predictions.
The Improved Evidence Lower Bound is given by:
$L_{\text{IELBO}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\hat{x}_{T}^{0}\right\|^{2},$ (16)
where $x_{T}^{0}$ and $\hat{x}_{T}^{0}$ represent the true and predicted values of the time series, respectively. This formulation combines the strengths of both probabilistic diffusion models and deterministic linear models, leading to more accurate forecasts.
5 EXPERIMENTS
Table 1: Overview of the dataset characteristics. Here, $T$ denotes the target series used for prediction.
Dataset ETTh1 ETTm1 Wind Exchange Weather Electricity
Dimension 7 7 7 8 21 321
Frequency 1 hour 15 mins 15 mins 1 day 10 mins 1 hour
T(Steps) 168 192 192 14 672 168
Table 2: The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
Dataset T CDPM iTransformer Patchmixer TSMixer PatchTST DLinear Diffusion-TS D3VAE CSDI
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 168 0.4364 0.4363 0.4433 0.4429 0.4610 0.4560 0.4418 0.4401 0.4401 0.4463 0.4372 0.4420 1.6559 1.0400 0.9523 0.8618 1.1090 0.8009
ETTh2 168 0.3393 0.3826 0.4009 0.4156 0.3731 0.4055 0.3654 0.3981 0.3805 0.4036 0.3976 0.4233 3.2225 1.4736 5.3165 1.9796 2.0700 1.0748
ETTm1 192 0.3532 0.3781 0.3685 0.3916 0.3731 0.3983 0.3673 0.3924 0.3626 0.3890 0.3706 0.3968 1.6769 1.0326 0.6341 0.6833 1.0770 0.7788
ETTm2 192 0.2303 0.3009 0.2535 0.3184 0.2643 0.3226 0.2495 0.3108 0.2472 0.3129 0.2498 0.3224 2.9380 1.3732 4.1001 1.5884 1.6051 0.9397
Electricity 168 0.1606 0.2526 0.1531 0.2461 0.1559 0.2505 0.1563 0.2497 0.1537 0.2494 0.1628 0.2581 1.4888 0.9935 1.7061 0.9273 —— ——
Exchange 14 0.0204 0.0979 0.0719 0.1896 0.0430 0.1150 0.0209 0.0955 0.0293 0.1178 0.0412 0.1320 2.7192 1.3573 4.8311 1.9942 1.3023 0.8719
Weather 672 0.3292 0.3373 0.3676 0.3709 0.3295 0.3411 0.3316 0.3430 0.3299 0.3400 0.3398 0.3801 —— —— —— —— 0.4842 0.4498
Wind 192 1.1262 0.7534 1.0653 0.7227 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
5.1 Experimental Settings
Datasets.
We evaluate the effectiveness of our proposed framework on multiple real-world time series datasets from various domains. The datasets include ETTh and ETTm Zhou et al. (2021), which represent electricity transformer temperature data; Exchange, which contains exchange rates from eight countries; and Weather [^3] https://www.bgc-jena.mpg.de/wetter/, consisting of 21 meteorological indicators sampled at 10-minute intervals. Additionally, we use the Electricity dataset [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which records electricity consumption from 321 clients, and Wind Li et al. (2022), comprising wind power measurements sampled every 15 minutes from 2020 to 2021.
To ensure consistency in chronological data splitting, we adopt a 6:2:2 split for the ETTh and ETTm datasets, and a 7:1:2 split for the Wind, Weather, Electricity, and Exchange datasets. This setup enables comprehensive testing of the framework’s generalization across different temporal segments. Consistent with TimeDiff Shen and Kwok (2023), we evaluate performance using historical window lengths of ${96,192,720,1440}$ to cover varying forecasting horizons.
Baselines.
We compare our proposed Conditional Denoising Polynomial Modeling (CDPM) framework against a diverse set of baseline models.
  • Diffusion models: CSDI Tashiro et al. (2021), Diffusion-TS Yuan and Qiao (2024).
  • Transformer-based models: iTransformer Liu et al. (2023), PatchTST Nie et al. (2023).
  • Linear-based models: DLinear Zeng et al. (2023), TSMixer Ekambaram et al. (2023).
  • Hybrid model: PatchMixer Gong et al. (2023), D3VAE Li et al. (2022).
Implementation Details.
Our proposed CDPM is implemented using PyTorch and trained with Exponential Learning Rate Scheduler Paszke et al. (2019), starting with an initial learning rate of $1\times 10^{-3}$ and a batch size of 16. Early stopping with a patience of 10 epochs is employed to prevent overfitting. The diffusion process utilizes $K=50$ steps with a cosine variance schedule Rasul et al. (2021) from $\beta_{1}=10^{-4}$ to $\beta_{K}=0.5$. We adopt the DDIM sampler Song et al. (2021). All experiments are conducted on a workstation equipped with an NVIDIA GeForce RTX 4090 GPU.
5.2 Experiment Results
Main Results.
The experimental results in Table 2 strongly validate the effectiveness of the CDPM for time series forecasting. CDPM’s decoupling strategy, which separates trend and seasonal components, allows for more accurate modeling of temporal patterns. It performs particularly well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise. CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics. On high-dimensional datasets like Electricity and Weather, CDPM maintains superior accuracy. Unlike CSDI and Diffusion-TS, which struggle with long forecast horizons, CDPM handles long-term dependencies effectively. On volatile datasets like Wind, CDPM captures sharp fluctuations and achieves competitive results, though further refinement could improve performance on highly noisy data. Overall, CDPM shows strong generalization across various datasets, outperforming models like DLinear and Autoformer. In summary, CDPM’s decoupling strategy enhances performance, confirming the theoretical hypothesis that it improves accuracy and interpretability over traditional and diffusion-based models.
Ablation Study.
Table 3: Ablation study results on ETTh2 and ETTm2.
Method ETTh2 ETTm2
MSE MAE MSE MAE
CDPM 0.3393 0.3826 0.2303 0.3009
w/o CR 0.3404 0.3781 0.2320 0.3023
PLR 0.3433 0.3826 0.2450 0.3121
w/o INA 0.3853 0.4215 0.2459 0.3272
w/o CR+PLR 0.3442 0.3807 0.2335 0.3015
The ablation experiments highlight the importance of each component in the CDPM framework, especially for handling complex temporal dynamics in datasets with varying granularities like ETTh2 and ETTm2. We present the results of ablation experiments in Table 4 to evaluate the contributions of individual components within the CDPM framework.
  • w/o CR: Removes the conditional information.
  • PLR: Replaces the PTM with a simple linear layer.
  • w/o INA: Removes instance normalization alignment.
  • w/o CR+PLR: Removes the conditional information and replaces PTM with a linear layer.
Removing the condition information injection module reduces performance, with a larger impact on ETTm2, where external factors influence intricate temporal patterns. Replacing the PTM with a linear layer causes a significant performance drop, particularly on ETTm2, underscoring the need for non-linear trend modeling. PTM is crucial for capturing long-term trends and periodic behaviors. Eliminating instance normalization leads to the largest performance decline, especially on ETTm2, highlighting its role in stabilizing training for high-resolution datasets. On ETTh2, instance normalization slightly improves performance, indicating its lesser importance when trends and seasonality are clearer. Removing both the condition information injection module and PTM further degrades performance, emphasizing the necessity of both components.
Figure 3: Parameter Sensitivity Analysis of CDPM: Hidden Dimension (Left), Patch Length (Middle), and Predicted Series Length (Right).
Figure 4: (a) Visual comparison of CDSM (left) and PTM (right) in modeling seasonal components. (b) Visual comparison of CDSM (left) and PTM (right) in modeling trend components.
Decoupling Validation Study.
Table 4: Decoupling validation study results on ETTh2 and ETTm2.
Method ETTh2 ETTm2
MSE MAE MSE MAE
CDPM 0.3393 0.3826 0.2303 0.3009
w/o DECOUPLING 0.3839 0.4106 0.2518 0.3157
REV 0.3541 0.3832 0.2452 0.3122
The Decoupling Validation Study emphasizes the importance of using appropriate models to represent the distinct components of the time series, while also providing evidence for the effectiveness and validity of the decoupling strategy. We present the results of ablation experiments in Table 4 to evaluate the validation of our decoupling framework.
  • w/o DECOUPLING: This variant directly applies the CDSM to the original time series.
  • REV: In this variant, PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with CDPM outperforming all ablated models. The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling. On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability. Applying the CDSM directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of CDPM, highlighting the importance of proper module assignment. These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures. For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but CDPM consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
PTM and CDSM Modeling Analysis.
To compare the performance of the specially designed components, PTM and CDSM, in modeling the trend and seasonal components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168. The results highlight the effectiveness of tailored modeling strategies for each component and the benefits of decoupled modeling in improving forecasting accuracy. As shown in Figure 4 (a), PTM excels in capturing intrinsic and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series. In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality. For the seasonal component, characterized by complex fluctuations, CDSM outperforms PTM, as shown in Figure 4(b). CDSM effectively captures sharp seasonal changes, while PTM focuses on underlying regularities in the fluctuations, though it may sometimes miss rapid shifts. These results validate each module’s modeling capabilities for its designated component. By assigning each module to the component it best models, CDPM captures both distinct trends and complex fluctuations, enhancing forecasting accuracy, as seen in both trend and seasonal visualizations.
Hyper-parameter Sensitiviy Analysis.
To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon, on model performance, we performed a hyperparameter sensitivity analysis using the ETTh1 dataset. The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, but larger values (e.g., $d_{\text{model}}=512$) yield diminishing returns and potential overfitting, highlighting the trade-off between model capacity and efficiency. Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global temporal patterns, while larger values introduce redundancy and reduce efficiency. Finally, CDPM performs consistently across prediction horizons, with PTM capturing trends and CDSM adapting to fluctuations, ensuring reliable forecasting. Overall, the analysis confirms CDPM’s balance between complexity and scalability, making it suitable for diverse forecasting tasks.
Figure 5: Visualization of trend (left) and seasonall (right) components of joint modeling for the electricity dataset.
Combined Modeling Capability Analysis.
To evaluate the framework’s interpretability and modeling capability, we performed a combined performance analysis of the PTM and CDSM within the CDPM, as shown in Figure 5. The CDPM decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its ability to model complex temporal dynamics. In Figure 5 (left), applied to the Electricity dataset with distinct periodic and trend characteristics, PTM captures intrinsic trends, with predictions closely matching actual trends. Figure 5 (right) highlights CDSM’s ability to model seasonal variations, with predicted components aligning well with actual patterns, compensating for PTM’s trend deviations and improving accuracy. The joint modeling of PTM and CDSM within the CDPM enhances both interpretability and predictive accuracy for datasets with trend and seasonal components.
6 Conclusion
In this paper, we proposed the Conditional Denoising Polynomial Modeling (CDPM) framework, a novel time series forecasting framework that separately models trend and seasonal components using specialized modules, trained in an end-to-end manner. The Polynomial Trend Module (PTM) captured intrinsic trends, while the Conditional Denoising Seasonal Module (CDSM) addressed complex seasonal fluctuations. By reformulating the Evidence Lower Bound, we were able to better integrate the training of both modules. This decoupling paradigm enhanced the modeling of temporal dynamics, resulting in improved forecasting accuracy. Extensive experiments on several real-world datasets validated the CDPM, demonstrating its consistent superiority over state-of-the-art models. The modular design, which handles distinct temporal patterns separately, advanced time series forecasting by improving both efficiency and accuracy.
Appendix
Appendix A Experiment Details
A.1 Denoising Network Details
The proposed Conditional Denoising Polynomial Modeling (CDPM) framework consists of several key subnetworks: the feature embedding network, the denoising decoder network, and the projector module. Table 5 details the input and output sizes of each subnetwork, where $d$ is the number of variables in the time series, $L$ is the sequence length, and $H$ is the hidden dimension of training process.
Table 5: Input and output sizes of the denoising networks.
Network Input size Output size
Embedding $T\times d$ $T\times d_{\text{model}}$
Denoising decoder $T\times d_{\text{model}}$ $T\times d_{\text{model}}$
Projector $T\times d_{\text{model}}$ $T\times d$
The key parameters are shown in Table 6. $d_{\text{hidden}}$ represents the hidden dimension used during the conditional information injection process, $d_{\text{model}}$ denotes the hidden dimension within the denoising network, $H$ indicates the patch size, and $LR$ refers to learning rate.
Table 6: Key parameters corresponding to each dataset.
Dataset $d_{\text{model}}$ $H$ $LR$
ETTh 256 8 0.001
ETTm 256 4 0.001
Exchange_rate 256 4 0.05
Weather 512 4 0.008
Wind 256 4 0.008
Electricity 1024 4 0.002
A.2 Datasets
We perform extensive experiments on several real-world time series datasets across diverse domains to assess the performance of our proposed framework. The datasets used in this study are: (i) ETTh and ETTm, comprising two years of electricity transformer temperature data from China, recorded at 1-hour and 15-minute intervals, respectively. These datasets offer valuable insights into various trends and complex variations in electricity usage. (ii) Exchange, which includes daily exchange rates from eight countries: Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore. This dataset presents challenges due to the complex dependencies and volatility inherent to financial time series. (iii) Weather [^5] https://www.bgc-jena.mpg.de/wetter/, containing 21 meteorological indicators recorded every 10 minutes from 2020 to 2021, providing data that captures seasonal patterns and sudden weather changes. (iv) Electricity [^6] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which provides hourly electricity consumption data from 321 clients over two years. This dataset is characterized by complex behaviors influenced by external factors. (v) Wind, containing wind power generation data recorded at 15-minute intervals from 2020 to 2021, reflecting dynamics shaped by daily routines and unforeseen events.
We split the ETTh and ETTm datasets into training, validation, and test sets using a 6:2:2 ratio, and employ a 7:1:2 split for the Weather, Electricity, and Exchange datasets. These splits ensure chronological consistency, crucial for realistic forecasting and preventing data leakage. Following TimeDiff, we select historical context lengths from ${96,192,720,1440}$, allowing for evaluation of model performance across various forecasting horizons.
A.3 Baselines
To comprehensively assess the effectiveness of our proposed CDPM, we compare it against a diverse set of baselines, each representing distinct methodological approaches:
A.4 Implementation Details
Our proposed CDPM model is implemented using PyTorch framework, providing a flexible and efficient platform for deep learning model development. The model is trained using an Exponential Learning Rate Scheduler, which dynamically adjusts learning rate throughout the training process to promote smoother convergence. Initially, learning rate is set to $1\times 10^{-3}$, and it decays exponentially over time. We also use a batch size of 16, which strikes a balance between training stability and computational efficiency. To prevent overfitting, early stopping is applied with a patience of 10 epochs, terminating training if validation performance does not improve over 10 consecutive epochs. The diffusion process in CDPM uses $K=50$ diffusion steps, following a cosine variance schedule, starting from $\beta_{1}=10^{-4}$ and increasing to $\beta_{K}=0.5$. This schedule balances the trade-off between noise addition and signal preservation during diffusion.
Several efficient learning-free samplers have been introduced to accelerate inference in diffusion models, such as DDIM, Analytic-DPM, and DPM-Solver. To expedite inference in our model, we utilize the Denoising Diffusion Implicit Models (DDIM) sampler, which reduces the number of diffusion steps required for sampling without compromising performance. This is particularly advantageous for real-time forecasting applications, where computational efficiency is paramount.
For model evaluation, we use the Mean Squared Error (MSE) to measure prediction accuracy. MSE is defined as:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2},$ (17)
where $\hat{y}_{i}$ is the predicted value, $y_{i}$ is the true value, and $n$ is the number of observations. Lower MSE values indicate better predictive performance. Additionally, the baseline loss is defined as:
$L_{\text{mse}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2},$ (18)
providing a reliable metric for optimization.
Algorithm 1 CDPM Training Algorithm.
1:  Require: Number of diffusion steps $K$, variance schedule $\{\beta_{k}\}_{k=1}^{K}$, historical series $x_{L}$, initialized model parameters $\theta$.
2:  repeat
3:     Normalize $x_{L}$ using its mean $\mu_{L}$ and variance $\sigma_{L}^{2}$, obtaining $x_{L}^{0}$; apply the same normalization to $x_{T}$, yielding $x_{T}^{0}$.
4:     Decompose $x_{L}^{0}$ into trend $x_{L,t}^{0}$ and seasonal $x_{L,s}^{0}$ components using (19) and decompose $x_{T}^{0}$ similarly using (20).
5:     Compute the cumulative product of the noise schedule: $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
6:     Sample a diffusion step $k$, and generate $x_{T,s}^{k}$ using (21).
7:     Predict the conditional information $\nabla_{\mathbf{x}_{T,s}}^{k}$ via (22).
8:     Use the predicted conditional information to estimate the seasonal component via (23).
9:     Predict the trend component according to (24).
10:     Integrate the predicted components using (25): $\hat{x}_{T}^{0}$.
11:     Denormalize $\hat{x}_{T}^{0}$ using $\mu_{L}$ and $\sigma_{L}^{2}$, obtaining $\hat{x}_{T}$.
12:     Compute the loss $L_{\text{mse}}(\theta)$ as defined in (26).
13:     Update the model parameters by taking a gradient descent step on $\nabla_{\theta}L_{\text{mse}}$.
14:  until convergence.
To account for the stochastic nature of diffusion processes, we average prediction results over 10 iterations for each time series model. This approach provides an effective estimate of the model’s performance by mitigating random fluctuations in predictions.
The experiments are conducted on a workstation with an NVIDIA GeForce RTX 4090 GPU (24 GB memory). This hardware setup ensures that computational resources do not limit model performance.
Algorithm 2 CDPM Inference Algorithm.
1:  Require: Model parameters $\theta$, number of diffusion steps $K$, variance schedule $\{\beta_{k}\}_{k=1}^{K}$, historical series $x_{L}$.
2:  Normalize $x_{L}$ using its mean $\mu_{L}$ and variance $\sigma_{L}^{2}$, yielding $x_{L}^{0}$, then decompose by (19) and (20).
3:  Initialize $x_{T,s}^{K}\sim\mathcal{N}(0,I)$.
4:  for $k=K$ to $1$ do
5:     Sample noise $\epsilon\sim\mathcal{N}(0,I)$ if $k>1$; otherwise, set $\epsilon=0$.
6:     Compute the cumulative product: $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
7:     Predict the conditional information $\nabla_{\mathbf{x}_{T,s}}^{k}$ using (22).
8:     Predict the seasonal component using formula (23).
9:     Update the denoised estimate:
$\hat{x}_{T,s}^{k-1}=\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}x_{T,s}^{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{x}_{T,s}^{0}+\sigma_{k}\epsilon.$
10:  end for
11:  Predict the trend component using formula (24).
12:  Integrate the trend and seasonal components, denormalize, and obtain the final output $\hat{x}_{T}$.
13:  Return $\hat{x}_{T}$.
Table 7: The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined. CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
Dataset L T CDPM iTransformer Patchmixer TSMixer PatchTST DLinear Diffusion-TS D3VAE CSDI
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 168 0.4329 0.4247 0.4324 0.4307 0.4337 0.4309 0.4224 0.4182 0.4300 0.4304 0.4325 0.4326 1.3530 0.8808 0.3531 0.5006 1.0454 0.7778
192 168 0.4313 0.4275 0.4292 0.4304 0.4347 0.4317 0.4281 0.4238 0.4217 0.4285 0.4257 0.4297 1.5407 1.0065 1.2732 1.0603 1.1051 0.7924
720 168 0.4424 0.4452 0.4750 0.4595 0.4534 0.4598 0.4340 0.4425 0.4675 0.4656 0.4259 0.4389 2.0399 1.1829 0.9371 0.8819 1.2069 0.8604
1440 168 0.4392 0.4478 0.4367 0.4512 0.5222 0.5015 0.4828 0.4759 0.4410 0.4607 0.4646 0.467 1.6900 1.0899 1.2457 1.0046 1.0785 0.7730
Average 0.4364 0.4363 0.4433 0.4429 0.4610 0.4560 0.4418 0.4401 0.4401 0.4463 0.4372 0.4420 1.6559 1.0400 0.9523 0.8618 1.1090 0.8009
ETTh2 96 168 0.3541 0.3837 0.3690 0.3912 0.3687 0.3925 0.3513 0.3816 0.3727 0.3904 0.3782 0.4076 3.3130 1.4568 2.5597 1.4723 2.1382 1.0701
192 168 0.3476 0.3869 0.3612 0.3931 0.3729 0.3987 0.3532 0.3859 0.3639 0.3923 0.3495 0.3863 3.9213 1.6013 6.0039 2.2822 1.5139 0.9363
720 168 0.3324 0.3811 0.4549 0.4445 0.3716 0.4128 0.3661 0.4012 0.3760 0.4012 0.3984 0.4288 2.0855 1.2501 6.7391 2.1875 1.4087 0.8958
1440 168 0.3231 0.3787 0.4185 0.4335 0.3792 0.4178 0.3911 0.4238 0.4093 0.4307 0.4645 0.4705 3.5703 1.5863 5.9632 1.9766 3.2192 1.3972
Average 0.3393 0.3826 0.4009 0.4156 0.3731 0.4055 0.3654 0.3981 0.3805 0.4036 0.3976 0.4233 3.2225 1.4736 5.3165 1.9796 2.0700 1.0748
ETTm1 96 192 0.3757 0.3846 0.3882 0.3941 0.3717 0.3913 0.3699 0.3857 0.3800 0.3926 0.3844 0.3920 1.6873 1.0646 0.5354 0.6012 0.9690 0.7273
192 192 0.3446 0.3715 0.3534 0.3812 0.3566 0.3857 0.3437 0.3753 0.3673 0.3889 0.3513 0.3822 1.2954 0.8786 0.5323 0.6288 1.1258 0.8013
720 192 0.3485 0.3774 0.3489 0.3864 0.3616 0.3908 0.3528 0.3875 0.3379 0.3787 0.3563 0.3888 2.1222 1.2104 0.8171 0.8229 1.0007 0.7421
1440 192 0.3438 0.3789 0.3837 0.4047 0.4025 0.4256 0.4029 0.4210 0.3653 0.3960 0.3904 0.4242 1.6025 0.9769 0.6516 0.6802 1.2127 0.8443
Average 0.3532 0.3781 0.3685 0.3916 0.3731 0.3983 0.3673 0.3924 0.3626 0.3890 0.3706 0.3968 1.6769 1.0326 0.6341 0.6833 1.0770 0.7788
ETTm2 96 192 0.2445 0.3043 0.2502 0.3084 0.2486 0.3099 0.2420 0.3019 0.2582 0.3201 0.2453 0.3133 3.4978 1.5438 0.5354 0.6012 1.4434 0.8721
192 192 0.2296 0.2959 0.2545 0.3146 0.2579 0.3146 0.2393 0.2988 0.2461 0.3089 0.2451 0.3094 2.3569 1.2538 1.4877 0.9955 1.5398 0.9198
720 192 0.2200 0.2984 0.2451 0.3174 0.2962 0.3344 0.2702 0.3262 0.2353 0.3071 0.2350 0.3137 2.7441 1.2724 2.6198 1.4873 1.6616 0.9615
1440 192 0.2269 0.3048 0.2644 0.3332 0.2545 0.3315 0.2467 0.3163 0.2492 0.3157 0.2738 0.3531 3.1531 1.4227 ##### 3.2698 1.7756 1.0054
Average 0.2303 0.3009 0.2535 0.3184 0.2643 0.3226 0.2495 0.3108 0.2472 0.3129 0.2498 0.3224 2.9380 1.3732 4.1001 1.5884 1.6051 0.9397
Electricity 96 168 0.1883 0.2730 0.1577 0.2515 0.1665 0.2567 0.1748 0.2625 0.1679 0.2592 0.1893 0.2763 1.4058 0.9506 0.6377 0.6256 —— ——
192 168 0.1612 0.2533 0.1566 0.2489 0.1503 0.2425 0.1536 0.2452 0.1511 0.2440 0.1622 0.2567 1.5622 1.0112 1.0692 0.8220 —— ——
720 168 0.1466 0.2414 0.1509 0.2428 0.1517 0.2481 0.1481 0.2464 0.1491 0.2479 0.1501 0.2493 1.4369 1.0072 4.0640 1.4426 —— ——
1440 168 0.1463 0.2428 0.1472 0.2412 0.1553 0.2546 0.1486 0.2446 0.1466 0.2465 0.1495 0.2501 1.5501 1.0050 1.0536 0.8189 —— ——
Average 0.1606 0.2526 0.1531 0.2461 0.1559 0.2505 0.1563 0.2497 0.1537 0.2494 0.1628 0.2581 1.4888 0.9935 1.7061 0.9273 —— ——
Exchange 96 14 0.0162 0.0830 0.0338 0.1331 0.0169 0.0834 0.0158 0.0807 0.0156 0.0825 0.0156 0.0799 2.5775 1.3205 3.1994 1.4322 1.1299 0.7540
192 14 0.0180 0.0892 0.1413 0.2728 0.0182 0.0885 0.0174 0.0849 0.0171 0.0883 0.0166 0.0838 3.1525 1.4267 7.1367 2.9513 1.1609 0.7874
720 14 0.0204 0.0992 0.0370 0.1441 0.0206 0.0999 0.0205 0.0964 0.0521 0.1683 0.0236 0.1092 2.2854 1.2949 1.0775 0.8829 1.4166 0.9583
1440 14 0.0269 0.1202 0.0753 0.2084 0.1166 0.1882 0.0300 0.1203 0.0324 0.1320 0.1091 0.2550 2.8615 1.3872 7.9106 2.7104 1.5019 0.9877
Average 0.0204 0.0979 0.0719 0.1896 0.0430 0.1150 0.0209 0.0955 0.0293 0.1178 0.0412 0.1320 2.7192 1.3573 4.8311 1.9942 1.3023 0.8719
Weather 96 672 0.3529 0.3470 0.3691 0.3626 0.3432 0.3420 0.3383 0.3384 0.3553 0.3490 0.3585 0.3988 —— —— —— —— 0.4796 0.4392
192 672 0.3401 0.3403 0.3602 0.3599 0.3285 0.3451 0.3245 0.3335 0.3398 0.3409 0.3502 0.3962 —— —— —— —— 0.5108 0.4526
720 672 0.3142 0.3314 0.3630 0.3732 0.3337 0.3426 0.3200 0.3404 0.3179 0.3383 0.3266 0.3659 —— —— —— —— 0.4697 0.4512
1440 672 0.3097 0.3304 0.3780 0.3877 0.3126 0.3349 0.3437 0.3596 0.3067 0.3318 0.3237 0.3595 —— —— —— —— 0.4765 0.4561
Average 0.3292 0.3373 0.3676 0.3709 0.3295 0.3411 0.3316 0.3430 0.3299 0.3400 0.3398 0.3801 —— —— —— —— 0.4842 0.4498
Wind 96 192 1.1953 0.7605 1.0926 0.7193 1.1628 0.7446 1.1471 0.7426 1.1796 0.7621 1.0439 0.7215 2.1130 1.0870 3.7312 1.3789 1.3909 0.8543
192 192 1.1693 0.7601 1.1040 0.7268 1.1904 0.7582 1.1597 0.7477 1.1637 0.7639 1.0481 0.7211 1.8822 1.0471 3.7236 1.4758 1.4510 0.8885
720 192 1.0969 0.7605 1.0309 0.7213 1.1576 0.7643 1.2393 0.7897 1.0970 0.7569 1.1007 0.7498 1.9993 1.1442 3.8525 1.4478 1.5825 0.9126
1440 192 1.0434 0.7326 1.0339 0.7234 1.0880 0.7658 1.4168 0.8580 1.1244 0.7597 1.1375 0.7587 2.2880 1.1957 3.4273 1.3756 1.5407 0.9212
Average 1.1262 0.7534 1.0653 0.7227 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
Appendix B Training and Inference Algorithm
Algorithms 1 and 2 illustrate the training algorithm and inference algorithm.
$X_{L,s}^{0}=X_{L}^{0}-X_{L,t}^{0},$ (19)
where the seasonal component $X_{L,s}^{0}$ is the residual after subtracting the trend.
The denoising process starts by embedding $X_{T,s}^{k}$:
$\mathbf{h}=\text{Emb}(X_{T,s}^{k}),$ (20)
where $\text{Emb}(\cdot)$ is a convolutional MLP that maps the noisy sequence to a hidden dimension while capturing local temporal features. Sinusoidal positional encodings for diffusion step $t$ are integrated via Adaptive Layer Normalization (AdaLN):
$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t}.$ (21)
where $\alpha_{t}$ and $\beta_{t}$ are learnable parameters projected from the diffusion step encoding.
The predicted statistics are used in the denoising process via Gaussian sampling:
$\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{T,i}+\hat{\sigma}_{T,i}\cdot z,$ (22)
where $z\sim\mathcal{N}(0,I)$. This enables the model to incorporate stochastic variability while aligning with historical seasonal patterns.
The denoised output $\hat{x}_{T,s}^{0}$ is combined with the conditional information through a weighted summation:
$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1}.$ (23)
where $\rho_{1}$ and $\rho_{2}$ are learnable parameters that balance the denoised series and conditional information, ensuring effective seasonal modeling while maintaining statistical properties.
The outputs from both pathways are combined via weighted summation to capture a wider range of trends:
$f_{\theta}(x_{T,t}^{0})=\lambda_{1}T_{\text{origin}}(x_{T,t}^{0})+\lambda_{2}T_{\text{sqrt}}(x_{T,t}^{0}),$ (24)
where $\lambda_{1}$ and $\lambda_{2}$ are learnable parameters that balance the contributions of the linear and root-based components.
The final prediction $\hat{x}_{T}^{0}$ calculated by integrating the trend and seasonal components as follows:
$\hat{x}_{T}^{0}=X_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{s,t}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\mid\nabla_{\mathbf{x}_{T,s}}^{k}\right)+f_{\theta}(x_{T,t}^{0}),$ (25)
where $X_{\theta}$ represents CDSM and $f_{\theta}$ corresponds to PTM.
The loss function in CDPM is derived from the ELBO:
$L_{\text{ELBO}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\hat{x}_{T}^{0}\right\|^{2},$ (26)
where $x_{T}$ and $\hat{x}_{T}$ represent the true value and the predicted value of the prediction window, respectively.
Table 8: Performance comparison across different variants on ETTh2 and ETTm2.
Dataset L T CDPM w/o CR PLR w/o INA w/o CR+PLR w/o DECOUPLING REV
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2 96 192 0.3541 0.3837 0.3690 0.3889 0.3518 0.3819 0.3650 0.3873 0.3764 0.3940 0.3609 0.3874 0.3712 0.3947
192 192 0.3476 0.3869 0.3480 0.3801 0.3466 0.3822 0.3403 0.3811 0.3426 0.3745 0.4039 0.4174 0.3878 0.4069
720 192 0.3324 0.3811 0.3287 0.3732 0.3395 0.3806 0.3314 0.3756 0.3444 0.3782 0.4015 0.4258 0.3904 0.4153
1440 192 0.3231 0.3787 0.3158 0.3701 0.3354 0.3856 0.3254 0.3819 0.3529 0.3862 0.3691 0.4118 0.3801 0.4139
Average 0.3393 0.3826 0.3404 0.3781 0.3433 0.3826 0.3405 0.3815 0.3541 0.3832 0.3839 0.4106 0.3824 0.4077
ETTm2 96 192 0.2445 0.3043 0.2461 0.3042 0.2453 0.3049 0.2465 0.3043 0.2585 0.3370 0.2435 0.3039 0.2456 0.3057
192 192 0.2296 0.2959 0.2307 0.2953 0.2307 0.2970 0.2297 0.2961 0.2642 0.3417 0.2396 0.3013 0.2502 0.3102
720 192 0.2200 0.2984 0.2279 0.3036 0.2379 0.3108 0.2192 0.2986 0.2331 0.3146 0.2488 0.3159 0.2517 0.3246
1440 192 0.2269 0.3048 0.2232 0.3061 0.2659 0.3356 0.2268 0.3049 0.2295 0.3172 0.2632 0.3383 0.2737 0.3418
Average 0.2303 0.3009 0.2320 0.3023 0.2450 0.3121 0.2305 0.3010 0.2463 0.3276 0.2488 0.3148 0.2553 0.3206
Appendix C DERIVED ELBO for CDPM
We begin by considering the standard diffusion model, which involves a forward process where data is gradually corrupted by noise, and a backward process to denoise it.
In the forward diffusion process, the input $x^{0}$ is transformed into a noisy vector $x^{k}$ through the following steps:
$q(x^{k}|x^{k-1})$ $=\mathcal{N}\left(x^{k};\sqrt{1-\beta_{k}}\,x^{k-1},\beta_{k}\,I\right),$ (27)
where $\beta_{k}$ is the variance schedule, and $I$ is the identity matrix.
By composing these steps, we can express the distribution of $x^{k}$ given $x^{0}$ as:
$q(x^{k}|x^{0})$ $=\mathcal{N}\left(x^{k};\sqrt{\bar{\alpha}_{k}}\,x^{0},\left(1-\bar{\alpha}_{k}\right)I\right),$ (28)
where $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
Thus, $x^{k}$ can be directly expressed as:
$x^{k}$ $=\sqrt{\bar{\alpha}_{k}}\,x^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (29)
The backward denoising process is defined by a neural network that predicts the mean and variance:
$p_{\theta}(x^{k-1}|x^{k})$ $=\mathcal{N}\left(x^{k-1};\mu_{\theta}(x^{k},k),\Sigma_{\theta}(x^{k},k)\right),$ (30)
where $\Sigma_{\theta}(x^{k},k)$ is fixed as $\sigma_{k}^{2}I$, and $\mu_{\theta}(x^{k},k)$ is modeled by a neural network.
The training objective minimizes the KL divergence between the forward and reverse processes:
$L_{k}$ $=D_{\text{KL}}\left(q(x^{k-1}|x^{k},x^{0})\,\|\,p_{\theta}(x^{k-1}|x^{k})\right).$ (31)
For stable training, $q(x^{k-1}|x^{k},x^{0})$ is simplified as:
$q(x^{k-1}|x^{k},x^{0})$ $=\mathcal{N}\left(x^{k-1};\mu_{k}(x^{0},x^{k}),\tilde{\beta}_{k}I\right),$ (32)
where $\mu_{k}(x^{0},x^{k})$ is expressed as:
$\mu_{k}(x^{0},x^{k})$ $=\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\,x^{0}+\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\,x^{k}.$ (33)
The training objective simplifies to:
$L_{k}$ $=\frac{1}{2\sigma_{k}^{2}}\,\mathbb{E}_{k}\left[\left\|\mu_{k}(x^{0},x^{k})-\mu_{\theta}(x^{k},k)\right\|^{2}\right].$ (34)
The mean of $x^{k-1}$, denoted $\mu_{x}(x_{\theta})$, is calculated as:
$\mu_{x}(x_{\theta})$ $=\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\,x^{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\,x_{\theta}(x^{k},k).$ (35)
The corresponding simplified loss is:
$L_{x}$ $=\omega\mathbb{E}_{k,x^{0},\epsilon}\left[\left\|x^{0}-x_{\theta}(x^{k},k)\right\|^{2}\right],$ (36)
where the parameter $\omega$ is a constant, and during training, it is typically set to 1.
Now , we introduce normalize the $x_{L}$ and $x_{T}$, and then decompose the condition window $x_{L}^{0}$ and predicted series $x_{T}^{0}$ into its trend and seasonal components:
$x_{L}^{0}$ $=x_{L,t}^{0}+x_{L,s}^{0}.$ (37)
$x_{T}^{0}$ $=x_{T,t}^{0}+x_{T,s}^{0}.$ (38)
We predict the trend component $x_{t}^{0}$ using $f_{\theta}$:
$\hat{x}_{T,t}^{0}$ $=f_{\theta}(x_{L,t}^{0}).$ (39)
Next, the seasonal component $x_{s}^{0}$ undergoes the diffusion process. The noisy version of the seasonal component, denoted by $r_{s}^{k}$, is obtained by adding noise to $x_{s}^{0}$:
$x_{T,s}^{k}$ $=\sqrt{\bar{\alpha}_{k}}x_{T,s}^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon.$ (40)
CDPM iTransformer PatchMixer TSMixer PatchTST DLinear
Figure 6: Comparison of CDPM and deterministic models on ETTh1 Dataset.
CDPM iTransformer PatchMixer TSMixer PatchTST DLinear
Figure 7: Comparison of CDPM and deterministic models on ETTh2 Dataset.
We improve the denoising process by integrating local statistical features derived from the historical sequence $x_{L,s}^{0}$. The sequence is segmented into patches $P$, each with a length $H$, to capture complex seasonal variations. For each patch $P_{i}$, we compute the mean $\mu_{i}$ and variance $\sigma_{i}^{2}$, which encapsulate the local statistical properties:
$\mu_{i}=\frac{1}{H}\sum_{h=1}^{H}x_{L,s,i,h}.$ (41)
$\quad\sigma_{i}^{2}=\frac{1}{H}\sum_{h=1}^{H}\left(x_{L,s,i,h}-\mu_{i}\right)^{2}.$ (42)
These computed statistics are utilized by a multi-layer perceptron (MLP) to predict the mean $\hat{\mu}_{i+1}$ and variance $\hat{\sigma}_{i+1}^{2}$ for the next patch:
$\hat{\mu}_{i+1}=\text{MLP}(\mu_{i}).$ (43)
$\quad\hat{\sigma}_{i+1}^{2}=\text{MLP}(\sigma_{i}^{2}).$ (44)
The conditional input for denoising is then sampled as:
$\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{i+1}+\hat{\sigma}_{i+1}\cdot z,\quad z\sim\mathcal{N}(0,1).$ (45)
This allows the seasonal component $\hat{x}_{T,s}^{0}$ to be predicted as:
$\hat{x}_{T,s}^{0}=X_{\theta}(x_{T,s}^{k},k\mid\nabla_{\mathbf{x}_{T,s}}^{k}).$ (46)
The final reconstruction of the time series is obtained by combining the predicted trend and seasonal components:
$\hat{x}_{T}^{0}$ $=\hat{x}_{T,t}^{0}+\hat{x}_{T,s}^{0}.$ (47)
To train the model, we minimize the mean squared error (MSE) between the true and predicted values for both components. The individual losses are defined as follows:
The trend loss measures the error between the true and predicted trend components:
$L_{\text{trend}}$ $=\mathbb{E}_{x_{T}^{0}}\left[\left\|x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right\|^{2}\right].$ (48)
The seasonal loss captures the error between the true and predicted seasonal components:
$L_{\text{seasonal}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left[\left\|x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right\|^{2}\right].$ (49)
We introduce a cross-term loss to account for the interaction between the trend and seasonal components. This term ensures the model considers any dependency between these components during prediction.
The cross-term loss is defined as:
$L_{\text{cross}}$ $=2\mathbb{E}_{x_{T}^{0}}\left[\left(x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right)\cdot\mathbb{E}_{\epsilon}\left[\left(x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right)\mid x_{T}^{0}\right]\right]$ (50)
$=2\mathbb{E}_{x_{T}^{0},\epsilon}\left[\left(x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right)\cdot\left(x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right)\right].$ (51)
The total loss function is the combination of the trend, seasonal, and cross-term losses:
$L_{\text{mse}}$ $=L_{\text{trend}}+L_{\text{seasonal}}+L_{\text{cross}}.$ (52)
Expanding the total loss, we obtain:
$L_{\text{mse}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\left(\hat{x}_{T,t}^{0}+\hat{x}_{T,s}^{0}\right)\right\|^{2}.$ (53)
Finally, we denormalize and integrate the diffusion model into the seasonal component, the loss function becomes:
$L_{\text{mse}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2}.$ (54)
This total loss function balances the contributions of the trend and seasonal components while accounting for their interactions. By minimizing this loss, the model learns to reconstruct the time series with greater accuracy.
CDPM CSDI Diffusion-TS
Figure 8: Comparison of CDPM and DDPM Models on ETTh1 Dataset.
CDPM CSDI Diffusion-TS
Figure 9: Comparison of CDPM and DDPM Models on ETTh2 Dataset.
Appendix D EXPERIMENTAL RESULTS
D.1 Overall Performance Results
The experimental results presented in Table 7 provide compelling evidence of the effectiveness of the Conditional Denoising Polynomial Modeling (CDPM) framework for time series forecasting. This advantage stems from CDPM’s decoupling strategy, which separates trend and seasonal components for more precise modeling of temporal patterns. This validates the framework’s theoretical contributions, where decoupling was hypothesized to improve accuracy and interpretability, now confirmed by the experimental results.
D.2 Overall Ablation Experiments
We conduct a comprehensive ablation study to evaluate the impact of key components in the proposed CDPM architecture. Specifically, we examine the role of the condition information injection module, the PTM, and instance normalization in handling the complex temporal dynamics present in time series data. The study is performed on the ETTh2 and ETTm2 datasets, and the results are presented in Table 8.
D.3 Case Study
As shown in Fig.6 and Fig.7, we compare CDPM with deterministic and diffusion-based models on the ETTh1 and ETTh2 datasets. With a history length of 192, CDPM consistently outperforms other models in capturing trends and seasonal variations. On ETTh1, CDPM closely aligns with the ground truth, excelling in modeling complex seasonal fluctuations and trend shifts. In contrast, deterministic models like PatchMixer and TSMixer fail to capture rapid trend changes and local seasonal variations, while Autoformer and DLinear produce overly smooth or delayed predictions. CDPM strikes an optimal balance in handling complex time series dynamics. In Fig.8 and Fig.9, CDPM is compared with diffusion-based models like CSDI and Diffusion-TS. CDPM offers smoother, more accurate forecasts, while Diffusion-TS shows higher variability and lower reliability, especially on ETTh2. By avoiding over-smoothing, CDPM better handles complex temporal structures. These results underscore CDPM’s ability to capture both trends and fluctuations, outperforming deterministic models and delivering more reliable forecasts than diffusion-based models. The advantages of CDPM’s decoupling strategy and specialized modules are clear.

### v6

@@
 Time series forecasting models are becoming increasingly prevalent due to their critical role in decision-making across various domains.
-However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between its specific components.
+However, most existing approaches represent the coupled temporal patterns, often neglecting the distinction between their specific components.
 In particular, fluctuating patterns and smooth trends within time series exhibit distinct characteristics.
@@
 Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models.
-[^2] Our code is open-sourced and available at https://github.com/zjt-gpu/FDF.
+[^2] Our code is available at https://github.com/zjt-gpu/CDPM.
 1 Introduction
 Time series forecasting is a fundamental task across various domains, playing a crucial role in driving informed decision-making in industries such as finance, healthcare, and manufacturing Kashpruk et al.
+(2023).
+It involves analyzing historical patterns to model the dynamic evolution of underlying processes and predict future trends Cheng et al.
+(2025a).
+Accurate time series forecasting is essential for tasks such as demand prediction, resource allocation, enabling organizations to anticipate future needs and optimize operations Kilcioglu et al.
 (2017); Liu et al.
 (2024).
-It involves analyzing historical patterns to model the dynamic evolution of underlying processes and predict future trends Montgomery et al.
-(2015).
-Accurate time series forecasting is essential for tasks such as demand prediction, resource allocation, enabling organizations to anticipate future needs and optimize operations.
-Trastour et al.
-(2002); Kilcioglu et al.
-(2017); Drutsa et al.
-(2015)
 Figure 1: The forward diffusion process iteratively adds noise to the original time series, which causes the trend components to be overshadowed as the noise increasingly dominates.
-This makes it challenging to recover trend during the reverse denoising process.
-Over the years, various approaches have been proposed to address this challenge, ranging from classical statistical models, such as Autoregressive Integrated Moving Average (ARIMA) Asteriou and Hall (2011), to machine learning techniques Masini et al.
+Over the years, various approaches have been proposed to address this challenge Cheng et al.
+(2025b), ranging from classical statistical models, such as Autoregressive Integrated Moving Average (ARIMA) Asteriou and Hall (2011), to machine learning techniques Masini et al.
 (2023), and advanced deep learning models, including recurrent neural networks (RNNs) Lin et al.
@@
 (2021).
-Large Language Models (LLM) like Time-LLM and LLMTIME Jin et al.
+Large language models (LLMs) such as TimeLLM and LLMTIME Jin et al.
 (2023); Gruver et al.
@@
 (2021); Shen and Kwok (2023), becoming the leading paradigm for probabilistic models originally developed for generation, have proven effective in time series forecasting due to their ability to capture intricate stochastic processes.
-Diffusion-based methods have shown significant potential in time series forecasting.
-We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fail to account for the distinct dynamics of different patterns.
+Diffusion-based methods Lee et al.
+(2024) have shown significant potential in time series forecasting.
+We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fails to account for the distinct dynamics of different patterns.
 Specifically, noise is uniformly added across the entire time series, and denoising is then applied indiscriminately, treating all temporal components equally.
-However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require distinct treatment.
+However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require different treatment.
 Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties.
-A uniform noise addition and denoising strategy fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting.
-As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing forecasting accuracy.
-Employing suitable models to capture the different components is crucial.
-Single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s forecasting ability.
-Diffusion models, with their advantages in modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series.
+This strategy of uniform noise addition and denoising fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting.
+As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing accuracy.
+Typically, a single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s overall forecasting ability.
+Diffusion models, which excel at modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series.
 In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends.
-Therefore, by leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more comprehensively capture the dynamic variations within the series and enhance forecasting accuracy.
+Therefore, by strategically leveraging the strengths of both model types and tailoring them to the distinct characteristics of time series, we can more fully capture the dynamic variations and enhance the final forecasting results.
 To overcome these challenges, we propose a novel Conditional Denoising Polynomial Modeling (CDPM) framework that leverages the inductive bias of time series decomposition.
-By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to each component’s characteristics.
-Specifically, we employ carefully designed Conditional Denoising Seasonal Module (CDSM) to capture the high-frequency seasonal component, introducing an innovative denoising strategy conditioned on statistical properties derived from historical window.
-Concurrently, we model the intrinsic trend component using Polynomial Trend Module (PTM), which are adept at capturing gradual changes over time.
-Additionally, we reformulate the Improved Evidence Lower Bound (IELBO) to facilitate the joint training of the CDSM and PTM.
-This decoupling method enhances the framework’s ability to capture temporal dependencies and complex patterns.
+By separating the time series into trend and seasonal components, we enable decoupled modeling tailored to the characteristics of each component.
+Specifically, we employ a carefully designed Conditional Denoising Seasonal Module (CDSM) to capture the high-frequency seasonal component, introducing an innovative denoising strategy conditioned on statistical properties derived from historical window.
+Currently, we model the intrinsic trend component using the polynomial trend module (PTM), which is adept at capturing gradual changes.
+Additionally, we reformulate the Improved Evidence Lower Bound (ELBO) to facilitate joint training of the CDSM and PTM.
+This decoupling method of our framework improves the ability to capture temporal dependencies and complex patterns.
 The key contributions of this work are:
@@
 (2018) and SegRNN Lin et al.
-(2023), became popular for sequential dependencies, while convolutional neural networks (CNNs) like TCN Bai et al.
+(2023), became popular for sequential dependencies, while convolutional neural networks (CNNs) Cheng et al.
+(2024) like TCN Bai et al.
 (2018) and SCINet Liu et al.
@@
 (2021) have utilized self-attention mechanisms to efficiently capture long-range dependencies, excelling in high-dimensional and long-horizon tasks.
-Surprisingly, simpler models like DLinear Zeng et al.
+Surprisingly, well-tuned linear models such as DLinear Zeng et al.
 (2023) and TSMixer Ekambaram et al.
-(2023) have shown that well-tuned linear layers, focusing on long-term trends, can outperform complex frameworks, highlighting their efficiency.
+(2023) have demonstrated strong performance by effectively capturing long-term trends, often outperforming more complex architectures and showcasing their efficiency.
 Despite advances, deterministic models often struggle with volatile fluctuations, limiting their ability to model complex time series dynamics.
-2.2 Probabilistic Forecasting
+2.2 Diffusion-Based Time Series Forecasting
 Probabilistic models offer uncertainty estimation and the ability to model temporal distributions.
@@
 (2023) applied diffusion to anomaly detection and interpolation.
-Diffusion models have also been successfully used in flood forecasting Shao et al.
-(2024), stock market prediction Daiya et al.
-(2024), and EV load forecasting Li et al.
-(2024).
 Despite the advancements in both deterministic and probabilistic forecasting approaches, existing research typically applies each paradigm separately to represent time series, with limited efforts to integrate them into a unified modeling framework.
 However, effectively capturing the diverse characteristics of time series, such as the smooth trend and fluctuating seasonal components, requires leveraging both deterministic and probabilistic models.
-3 Notations and Problem Statement
+3 Problem Statement
 Time series represent numerical sequence ordered chronologically, typically recorded at regular temporal intervals.
@@
 By handling trend and seasonal components independently, CDPM allows each module to focus on its respective dynamics, leading to enhanced overall performance.
-4.2 Preprocessing for CDPM
+4.2 Temporal Feature Structuring
 Instance Normalization.
@@
 This ensures alignment between the windows and preserves both the temporal structure and stability in the forecasting process.
-The subsequent denormalization step restores the forecasted values to their original distribution, thereby maintaining consistency with input data.
+The subsequent denormalization step restores the predicted values to their original scale, ensuring consistency with the input data distribution.
 Decomposition Operation.
@@
 4.3 Conditional Denoising Seasonal Module
-Modeling fluctuating seasonal components in time series is challenging due to their non-stationary and intricate patterns.
-We propose the CDSM, based on denoising diffusion model, which is a popular paradigm in probabilistic modeling.
+Modeling fluctuating seasonal components in time series is challenging due to their non-stationarity and intricate patterns.
+We propose CDSM, based on a denoising diffusion model, a popular paradigm in probabilistic modeling.
 By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
@@
 As $k$ increases, the seasonal component becomes more noise-dominated.
-Reversed Conditional Denoise.
+Reverse Conditional Denoising Process.
 The denoising process starts by embedding $X_{T,s}^{k}$:
@@
 Sinusoidal positional encodings for diffusion step $t$ are integrated via Adaptive Layer Normalization (AdaLN):
-$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t}.$ (5)
+$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t},$ (5)
 where $\alpha_{t}$ and $\beta_{t}$ are learnable parameters projected from the diffusion step encoding.
@@
 $\hat{\sigma}_{T,i}^{2}=\text{MLP}_{\sigma}(\sigma_{L,i}^{2}),$ (9)
-where $\text{MLP}_{\mu}(\cdot)$ and $\text{MLP}_{\sigma}(\cdot)$ capture the mean and variance of historical window features, respectively.
-The predicted statistics are used in the denoising process via Gaussian sampling:
+where $\text{MLP}_{\mu}(\cdot)$ and $\text{MLP}_{\sigma}(\cdot)$ capture the mean and variance of the historical window features, respectively.
+These predicted statistics are subsequently used in the denoising process via Gaussian sampling:
 $\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{T,i}+\hat{\sigma}_{T,i}\cdot z,$ (10)
@@
 Finally, the denoised output $\hat{x}_{T,s}^{0}$ is combined with the conditional information through a weighted summation:
-$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1}.$ (11)
+$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1},$ (11)
 where $\rho_{1}$ and $\rho_{2}$ are learnable parameters that balance the denoised series and conditional information, ensuring effective seasonal modeling while maintaining statistical properties.
@@
 This mechanism allows the model to adapt to both linear and sub-linear growth patterns, enhancing its ability to forecast complex time series with diverse dynamics.
-4.5 Reconstruction and Optimization
+4.5 Optimization
 Integrating trend and seasonal components improves forecasting accuracy by combining their strengths for more comprehensive predictions.
 Finally, the derived loss function enables joint optimization, allowing both components to be learned concurrently and effectively.
-Reconstruction.
+Component Reconstruction.
 During the training and inference phases, after the trend component $\hat{x}_{T,t}^{0}$ is predicted by the PTM and the seasonal component $\hat{x}_{T,s}^{0}$ is recovered by the CDSM.
-The final prediction $\hat{x}_{T}^{0}$ calculated by integrating the trend and seasonal components as follows:
+The final prediction $\hat{x}_{T}^{0}$ is calculated by integrating the trend and seasonal components as follows:
 $\hat{x}_{T}^{0}=X_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{s,t}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\mid\nabla_{\mathbf{x}_{T,s}}^{k}\right)+f_{\theta}(x_{T,t}^{0}),$ (15)
-where $X_{\theta}(\cdot)$ represents the CDSM and $f_{\theta}(\cdot)$ corresponds to the PTM.
-Improved Evidence Lower Bound Optimization.
+where $X_{\theta}(\cdot)$ represents the CDSM and $f_{\theta}(\cdot)$ corresponds to the PTM, respectively.
+Improved ELBO Optimization.
 The final prediction $\hat{x}_{T}^{0}$ is subsequently denormalized to $\hat{x}_{T}$ using the mean and variance derived from the historical window, ensuring that the forecast is consistent with the original distribution of the time series.
@@
 Through derivation, we find that by modeling the seasonal component with CDSM and the trend component with PTM, the results can be jointly optimized and trained to improve the overall performance of our framework, leading to more accurate predictions.
-The Improved Evidence Lower Bound is given by:
+The improved ELBO is given by:
 $L_{\text{IELBO}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\hat{x}_{T}^{0}\right\|^{2},$ (16)
@@
 5 Experiments
+5.1 Experimental Settings
 Table 1: Overview of the dataset characteristics. Here, $T$ denotes the target series used for prediction.
@@
 Wind 192 1.1262 0.7534 1.0653 0.7227 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
-5.1 Experimental Settings
 Datasets.
@@
 Main Results.
-The experimental results in Table 2 strongly validate the effectiveness of the CDPM for time series forecasting.
+The experimental results in Table 2 validate the effectiveness of the CDPM for time series forecasting.
 CDPM’s decoupling strategy, which separates trend and seasonal components, allows for more accurate modeling of temporal patterns.
-It performs particularly well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise.
+It performs well on challenging datasets like ETTh1 and ETTh2, where traditional models such as Diffusion-TS and D3VAE struggle with varying time resolutions and high noise.
 CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics.
@@
 On volatile datasets like Wind, CDPM captures sharp fluctuations and achieves competitive results, though further refinement could improve performance on highly noisy data.
-Overall, CDPM shows strong generalization across various datasets, outperforming models like DLinear and Autoformer.
+Overall, CDPM shows strong generalization across various datasets.
 In summary, CDPM’s decoupling strategy enhances performance, confirming the theoretical hypothesis that it improves accuracy and interpretability over traditional and diffusion-based models.
@@
 CDPM 0.3393 0.3826 0.2303 0.3009
-w/o CR 0.3404 0.3781 0.2320 0.3023
-PLR 0.3433 0.3826 0.2450 0.3121
-w/o INA 0.3853 0.4215 0.2459 0.3272
-w/o CR+PLR 0.3442 0.3807 0.2335 0.3015
+CDPM1 0.3404 0.3781 0.2320 0.3023
+CDPM2 0.3433 0.3826 0.2450 0.3121
+CDPM3 0.3853 0.4215 0.2459 0.3272
+CDPM4 0.3442 0.3807 0.2335 0.3015
 The ablation experiments highlight the importance of each component in the CDPM framework, especially for handling complex temporal dynamics in datasets with varying granularities like ETTh2 and ETTm2.
 We present the results of ablation experiments in Table 3 to evaluate the contributions of individual components within the CDPM framework.
-  • w/o CR: Removes the conditional information.
-  • PLR: Replaces the PTM with a simple linear layer.
-  • w/o INA: Removes instance normalization alignment.
-  • w/o CR+PLR: Removes the conditional information and replaces PTM with a linear layer.
+  • CDPM1: Removes the conditional information.
+  • CDPM2: Replaces the PTM with a simple linear layer.
+  • CDPM3: Removes instance normalization alignment.
+  • CDPM4: Removes the conditional information and replaces PTM with a linear layer.
 Removing the condition information injection module reduces performance, with a larger impact on ETTm2, where external factors influence intricate temporal patterns.
@@
 Removing both the condition information injection module and PTM further degrades performance, emphasizing the necessity of both components.
-Figure 3: Parameter Sensitivity Analysis of CDPM: Hidden Dimension (Left), Patch Length (Middle), and Predicted Series Length (Right).
+Figure 3: Parameter sensitivity analysis of CDPM: hidden dimension (left), patch length (middle), and predicted length (right).
 Figure 4: (a) Visual comparison of CDSM (left) and PTM (right) in modeling seasonal components. (b) Visual comparison of CDSM (left) and PTM (right) in modeling trend components.
-Decoupling Validation Study.
+Evaluation of the Decoupling Strategy.
+The Decoupling Validation Study emphasizes the importance of using appropriate models to represent the distinct components of the time series, while also providing evidence for the effectiveness and validity of the decoupling strategy.
+We present the results of ablation experiments in Table 4 to evaluate the validation of our decoupling framework.
+  • CDPM5: This variant directly applies the CDSM to the original time series.
+  • CDPM6: In this variant, PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
+The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with our framework CDPM outperforming all ablated models.
+The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling.
+On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability.
+Applying our framework directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of our framework, highlighting the importance of proper module assignment.
+These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures.
+For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but our framework consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
 Table 4: Decoupling validation study results on ETTh2 and ETTm2.
@@
 CDPM 0.3393 0.3826 0.2303 0.3009
-w/o DECOUPLING 0.3839 0.4106 0.2518 0.3157
-REV 0.3541 0.3832 0.2452 0.3122
-The Decoupling Validation Study emphasizes the importance of using appropriate models to represent the distinct components of the time series, while also providing evidence for the effectiveness and validity of the decoupling strategy.
-We present the results of ablation experiments in Table 4 to evaluate the validation of our decoupling framework.
-  • w/o DECOUPLING: This variant directly applies the CDSM to the original time series.
-  • REV: In this variant, PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
-The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with CDPM outperforming all ablated models.
-The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling.
-On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability.
-Applying the CDSM directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of CDPM, highlighting the importance of proper module assignment.
-These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures.
-For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but CDPM consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
+CDPM5 0.3839 0.4106 0.2518 0.3157
+CDPM6 0.3541 0.3832 0.2452 0.3122
 PTM and CDSM Modeling Analysis.
@@
 In contrast, applying CDSM to the trend component introduces unnecessary noise and overfitting to minor variations, diminishing trend quality.
-For the seasonal component, characterized by complex fluctuations, CDSM outperforms PTM, as shown in Figure 4(b).
+For the seasonal component, which exhibits complex fluctuations, the seasonal module demonstrates superior performance, as shown in Figure 4 (b).
 CDSM effectively captures sharp seasonal changes, while PTM focuses on underlying regularities in the fluctuations, though it may sometimes miss rapid shifts.
 These results validate each module’s modeling capabilities for its designated component.
-By assigning each module to the component it best models, CDPM captures both distinct trends and complex fluctuations, enhancing forecasting accuracy, as seen in both trend and seasonal visualizations.
-Hyper-parameter Sensitiviy Analysis.
-To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon, on model performance, we performed a hyperparameter sensitivity analysis using the ETTh1 dataset.
-The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, but larger values (e.g., $d_{\text{model}}=512$) yield diminishing returns and potential overfitting, highlighting the trade-off between model capacity and efficiency.
-Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global temporal patterns, while larger values introduce redundancy and reduce efficiency.
-Finally, CDPM performs consistently across prediction horizons, with PTM capturing trends and CDSM adapting to fluctuations, ensuring reliable forecasting.
-Overall, the analysis confirms CDPM’s balance between complexity and scalability, making it suitable for diverse forecasting tasks.
+Hyper-Parameter Sensitivity Analysis.
+To investigate the impact of key design choices in CDPM, such as hidden dimension size ($d_{\text{model}}$), patch length ($H$), and prediction horizon on model performance, we conducted a hyperparameter sensitivity analysis using the ETTh1 dataset.
+The results show that increasing $d_{\text{model}}$ initially improves performance, peaking at $d_{\text{model}}=256$, while larger values (e.g., $d_{\text{model}}=512$) offer diminishing returns and risk overfitting, highlighting the trade-off between model capacity and efficiency.
+Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global patterns, whereas larger values introduce redundancy and reduce efficiency.
+Finally, our framework performs consistently across prediction horizons, with the trend module capturing long-term structure and the seasonal module adapting to fluctuations, ensuring reliable forecasts.
+Overall, the analysis confirms the balance between complexity and scalability in our framework, making it suitable for diverse forecasting tasks.
 Figure 5: Visualization of trend (left) and seasonall (right) components of joint modeling for the electricity dataset.
 Combined Modeling Capability Analysis.
-To evaluate the framework’s interpretability and modeling capability, we performed a combined performance analysis of the PTM and CDSM within the CDPM, as shown in Figure 5.
-The CDPM decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its ability to model complex temporal dynamics.
-In Figure 5 (left), applied to the Electricity dataset with distinct periodic and trend characteristics, PTM captures intrinsic trends, with predictions closely matching actual trends.
-Figure 5 (right) highlights CDSM’s ability to model seasonal variations, with predicted components aligning well with actual patterns, compensating for PTM’s trend deviations and improving accuracy.
-The joint modeling of PTM and CDSM within the CDPM enhances both interpretability and predictive accuracy for datasets with trend and seasonal components.
+To assess the interpretability and predictive performance of the framework, we conducted a joint analysis of the PTM and CDSM modules on the Electricity dataset, as shown in Figure 5.
+The framework decouples multidimensional time series, capturing diverse trends and cyclic patterns, demonstrating its capacity to model complex temporal dynamics.
+In Figure 5 (left), applied to the Electricity dataset with clear periodic and trend structures, the trend module captures the underlying patterns, with predictions closely matching actual values.
+Figure 5 (right) illustrates the seasonal module’s ability to model periodic variations, where predicted components align well with the observed sequences, compensating for deviations in trend estimation and improving accuracy.
+The integration of deterministic and diffusion-based modules within the framework enhances both interpretability and predictive accuracy across datasets with combined components.
 6 Conclusion
-In this paper, we proposed the Conditional Denoising Polynomial Modeling (CDPM) framework, a novel time series forecasting framework that separately models trend and seasonal components using specialized modules, trained in an end-to-end manner.
+In this paper, we proposed the Conditional Denoising Polynomial Modeling (CDPM), a novel time series forecasting framework that separately models trend and seasonal components using specialized modules, trained in an end-to-end manner.
 The Polynomial Trend Module (PTM) captured intrinsic trends, while the Conditional Denoising Seasonal Module (CDSM) addressed complex seasonal fluctuations.
@@
 This decoupling paradigm enhanced the modeling of temporal dynamics, resulting in improved forecasting accuracy.
-Extensive experiments on several real-world datasets validated the CDPM, demonstrating its consistent superiority over state-of-the-art models.
-The modular design, which handles distinct temporal patterns separately, advanced time series forecasting by improving both efficiency and accuracy.
-Appendix
-Appendix A Experiment Details
-A.1 Denoising Network Details
-The proposed Conditional Denoising Polynomial Modeling (CDPM) framework consists of several key subnetworks: the feature embedding network, the denoising decoder network, and the projector module.
-Table 5 details the input and output sizes of each subnetwork, where $d$ is the number of variables in the time series, $L$ is the sequence length, and $H$ is the hidden dimension of training process.
-Table 5: Input and output sizes of the denoising networks.
-Network Input size Output size
-Embedding $T\times d$ $T\times d_{\text{model}}$
-Denoising decoder $T\times d_{\text{model}}$ $T\times d_{\text{model}}$
-Projector $T\times d_{\text{model}}$ $T\times d$
-The key parameters are shown in Table 6.
-$d_{\text{hidden}}$ represents the hidden dimension used during the conditional information injection process, $d_{\text{model}}$ denotes the hidden dimension within the denoising network, $H$ indicates the patch size, and $LR$ refers to learning rate.
-Table 6: Key parameters corresponding to each dataset.
-Dataset $d_{\text{model}}$ $H$ $LR$
-ETTh 256 8 0.001
-ETTm 256 4 0.001
-Exchange_rate 256 4 0.05
-Weather 512 4 0.008
-Wind 256 4 0.008
-Electricity 1024 4 0.002
-A.2 Datasets
-We perform extensive experiments on several real-world time series datasets across diverse domains to assess the performance of our proposed framework.
-The datasets used in this study are: (i) ETTh and ETTm, comprising two years of electricity transformer temperature data from China, recorded at 1-hour and 15-minute intervals, respectively.
-These datasets offer valuable insights into various trends and complex variations in electricity usage.
-(ii) Exchange, which includes daily exchange rates from eight countries: Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore.
-This dataset presents challenges due to the complex dependencies and volatility inherent to financial time series.
-(iii) Weather [^5] https://www.bgc-jena.mpg.de/wetter/, containing 21 meteorological indicators recorded every 10 minutes from 2020 to 2021, providing data that captures seasonal patterns and sudden weather changes.
-(iv) Electricity [^6] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, which provides hourly electricity consumption data from 321 clients over two years.
-This dataset is characterized by complex behaviors influenced by external factors.
-(v) Wind, containing wind power generation data recorded at 15-minute intervals from 2020 to 2021, reflecting dynamics shaped by daily routines and unforeseen events.
-We split the ETTh and ETTm datasets into training, validation, and test sets using a 6:2:2 ratio, and employ a 7:1:2 split for the Weather, Electricity, and Exchange datasets.
-These splits ensure chronological consistency, crucial for realistic forecasting and preventing data leakage.
-Following TimeDiff, we select historical context lengths from ${96,192,720,1440}$, allowing for evaluation of model performance across various forecasting horizons.
-A.3 Baselines
-To comprehensively assess the effectiveness of our proposed CDPM, we compare it against a diverse set of baselines, each representing distinct methodological approaches:
-A.4 Implementation Details
-Our proposed CDPM model is implemented using PyTorch framework, providing a flexible and efficient platform for deep learning model development.
-The model is trained using an Exponential Learning Rate Scheduler, which dynamically adjusts learning rate throughout the training process to promote smoother convergence.
-Initially, learning rate is set to $1\times 10^{-3}$, and it decays exponentially over time.
-We also use a batch size of 16, which strikes a balance between training stability and computational efficiency.
-To prevent overfitting, early stopping is applied with a patience of 10 epochs, terminating training if validation performance does not improve over 10 consecutive epochs.
-The diffusion process in CDPM uses $K=50$ diffusion steps, following a cosine variance schedule, starting from $\beta_{1}=10^{-4}$ and increasing to $\beta_{K}=0.5$.
-This schedule balances the trade-off between noise addition and signal preservation during diffusion.
-Several efficient learning-free samplers have been introduced to accelerate inference in diffusion models, such as DDIM, Analytic-DPM, and DPM-Solver.
-To expedite inference in our model, we utilize the Denoising Diffusion Implicit Models (DDIM) sampler, which reduces the number of diffusion steps required for sampling without compromising performance.
-This is particularly advantageous for real-time forecasting applications, where computational efficiency is paramount.
-For model evaluation, we use the Mean Squared Error (MSE) to measure prediction accuracy. MSE is defined as:
-$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}-y_{i})^{2},$ (17)
-where $\hat{y}_{i}$ is the predicted value, $y_{i}$ is the true value, and $n$ is the number of observations.
-Lower MSE values indicate better predictive performance.
-Additionally, the baseline loss is defined as:
-$L_{\text{mse}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2},$ (18)
-providing a reliable metric for optimization.
-Algorithm 1 CDPM Training Algorithm.
-1:  Require: Number of diffusion steps $K$, variance schedule $\{\beta_{k}\}_{k=1}^{K}$, historical series $x_{L}$, initialized model parameters $\theta$.
-2:  repeat
-3:     Normalize $x_{L}$ using its mean $\mu_{L}$ and variance $\sigma_{L}^{2}$, obtaining $x_{L}^{0}$; apply the same normalization to $x_{T}$, yielding $x_{T}^{0}$.
-4:     Decompose $x_{L}^{0}$ into trend $x_{L,t}^{0}$ and seasonal $x_{L,s}^{0}$ components using (19) and decompose $x_{T}^{0}$ similarly using (20).
-5:     Compute the cumulative product of the noise schedule: $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
-6:     Sample a diffusion step $k$, and generate $x_{T,s}^{k}$ using (21).
-7:     Predict the conditional information $\nabla_{\mathbf{x}_{T,s}}^{k}$ via (22).
-8:     Use the predicted conditional information to estimate the seasonal component via (23).
-9:     Predict the trend component according to (24).
-10:     Integrate the predicted components using (25): $\hat{x}_{T}^{0}$.
-11:     Denormalize $\hat{x}_{T}^{0}$ using $\mu_{L}$ and $\sigma_{L}^{2}$, obtaining $\hat{x}_{T}$.
-12:     Compute the loss $L_{\text{mse}}(\theta)$ as defined in (26).
-13:     Update the model parameters by taking a gradient descent step on $\nabla_{\theta}L_{\text{mse}}$.
-14:  until convergence.
-To account for the stochastic nature of diffusion processes, we average prediction results over 10 iterations for each time series model.
-This approach provides an effective estimate of the model’s performance by mitigating random fluctuations in predictions.
-The experiments are conducted on a workstation with an NVIDIA GeForce RTX 4090 GPU (24 GB memory). This hardware setup ensures that computational resources do not limit model performance.
-Algorithm 2 CDPM Inference Algorithm.
-1:  Require: Model parameters $\theta$, number of diffusion steps $K$, variance schedule $\{\beta_{k}\}_{k=1}^{K}$, historical series $x_{L}$.
-2:  Normalize $x_{L}$ using its mean $\mu_{L}$ and variance $\sigma_{L}^{2}$, yielding $x_{L}^{0}$, then decompose by (19) and (20).
-3:  Initialize $x_{T,s}^{K}\sim\mathcal{N}(0,I)$.
-4:  for $k=K$ to $1$ do
-5:     Sample noise $\epsilon\sim\mathcal{N}(0,I)$ if $k>1$; otherwise, set $\epsilon=0$.
-6:     Compute the cumulative product: $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
-7:     Predict the conditional information $\nabla_{\mathbf{x}_{T,s}}^{k}$ using (22).
-8:     Predict the seasonal component using formula (23).
-9:     Update the denoised estimate:
-$\hat{x}_{T,s}^{k-1}=\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}x_{T,s}^{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\hat{x}_{T,s}^{0}+\sigma_{k}\epsilon.$
-10:  end for
-11:  Predict the trend component using formula (24).
-12:  Integrate the trend and seasonal components, denormalize, and obtain the final output $\hat{x}_{T}$.
-13:  Return $\hat{x}_{T}$.
-Table 7: The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined.
-CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
-Dataset L T CDPM iTransformer Patchmixer TSMixer PatchTST DLinear Diffusion-TS D3VAE CSDI
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 96 168 0.4329 0.4247 0.4324 0.4307 0.4337 0.4309 0.4224 0.4182 0.4300 0.4304 0.4325 0.4326 1.3530 0.8808 0.3531 0.5006 1.0454 0.7778
-192 168 0.4313 0.4275 0.4292 0.4304 0.4347 0.4317 0.4281 0.4238 0.4217 0.4285 0.4257 0.4297 1.5407 1.0065 1.2732 1.0603 1.1051 0.7924
-720 168 0.4424 0.4452 0.4750 0.4595 0.4534 0.4598 0.4340 0.4425 0.4675 0.4656 0.4259 0.4389 2.0399 1.1829 0.9371 0.8819 1.2069 0.8604
-1440 168 0.4392 0.4478 0.4367 0.4512 0.5222 0.5015 0.4828 0.4759 0.4410 0.4607 0.4646 0.467 1.6900 1.0899 1.2457 1.0046 1.0785 0.7730
-Average 0.4364 0.4363 0.4433 0.4429 0.4610 0.4560 0.4418 0.4401 0.4401 0.4463 0.4372 0.4420 1.6559 1.0400 0.9523 0.8618 1.1090 0.8009
-ETTh2 96 168 0.3541 0.3837 0.3690 0.3912 0.3687 0.3925 0.3513 0.3816 0.3727 0.3904 0.3782 0.4076 3.3130 1.4568 2.5597 1.4723 2.1382 1.0701
-192 168 0.3476 0.3869 0.3612 0.3931 0.3729 0.3987 0.3532 0.3859 0.3639 0.3923 0.3495 0.3863 3.9213 1.6013 6.0039 2.2822 1.5139 0.9363
-720 168 0.3324 0.3811 0.4549 0.4445 0.3716 0.4128 0.3661 0.4012 0.3760 0.4012 0.3984 0.4288 2.0855 1.2501 6.7391 2.1875 1.4087 0.8958
-1440 168 0.3231 0.3787 0.4185 0.4335 0.3792 0.4178 0.3911 0.4238 0.4093 0.4307 0.4645 0.4705 3.5703 1.5863 5.9632 1.9766 3.2192 1.3972
-Average 0.3393 0.3826 0.4009 0.4156 0.3731 0.4055 0.3654 0.3981 0.3805 0.4036 0.3976 0.4233 3.2225 1.4736 5.3165 1.9796 2.0700 1.0748
-ETTm1 96 192 0.3757 0.3846 0.3882 0.3941 0.3717 0.3913 0.3699 0.3857 0.3800 0.3926 0.3844 0.3920 1.6873 1.0646 0.5354 0.6012 0.9690 0.7273
-192 192 0.3446 0.3715 0.3534 0.3812 0.3566 0.3857 0.3437 0.3753 0.3673 0.3889 0.3513 0.3822 1.2954 0.8786 0.5323 0.6288 1.1258 0.8013
-720 192 0.3485 0.3774 0.3489 0.3864 0.3616 0.3908 0.3528 0.3875 0.3379 0.3787 0.3563 0.3888 2.1222 1.2104 0.8171 0.8229 1.0007 0.7421
-1440 192 0.3438 0.3789 0.3837 0.4047 0.4025 0.4256 0.4029 0.4210 0.3653 0.3960 0.3904 0.4242 1.6025 0.9769 0.6516 0.6802 1.2127 0.8443
-Average 0.3532 0.3781 0.3685 0.3916 0.3731 0.3983 0.3673 0.3924 0.3626 0.3890 0.3706 0.3968 1.6769 1.0326 0.6341 0.6833 1.0770 0.7788
-ETTm2 96 192 0.2445 0.3043 0.2502 0.3084 0.2486 0.3099 0.2420 0.3019 0.2582 0.3201 0.2453 0.3133 3.4978 1.5438 0.5354 0.6012 1.4434 0.8721
-192 192 0.2296 0.2959 0.2545 0.3146 0.2579 0.3146 0.2393 0.2988 0.2461 0.3089 0.2451 0.3094 2.3569 1.2538 1.4877 0.9955 1.5398 0.9198
-720 192 0.2200 0.2984 0.2451 0.3174 0.2962 0.3344 0.2702 0.3262 0.2353 0.3071 0.2350 0.3137 2.7441 1.2724 2.6198 1.4873 1.6616 0.9615
-1440 192 0.2269 0.3048 0.2644 0.3332 0.2545 0.3315 0.2467 0.3163 0.2492 0.3157 0.2738 0.3531 3.1531 1.4227 ##### 3.2698 1.7756 1.0054
-Average 0.2303 0.3009 0.2535 0.3184 0.2643 0.3226 0.2495 0.3108 0.2472 0.3129 0.2498 0.3224 2.9380 1.3732 4.1001 1.5884 1.6051 0.9397
-Electricity 96 168 0.1883 0.2730 0.1577 0.2515 0.1665 0.2567 0.1748 0.2625 0.1679 0.2592 0.1893 0.2763 1.4058 0.9506 0.6377 0.6256 —— ——
-192 168 0.1612 0.2533 0.1566 0.2489 0.1503 0.2425 0.1536 0.2452 0.1511 0.2440 0.1622 0.2567 1.5622 1.0112 1.0692 0.8220 —— ——
-720 168 0.1466 0.2414 0.1509 0.2428 0.1517 0.2481 0.1481 0.2464 0.1491 0.2479 0.1501 0.2493 1.4369 1.0072 4.0640 1.4426 —— ——
-1440 168 0.1463 0.2428 0.1472 0.2412 0.1553 0.2546 0.1486 0.2446 0.1466 0.2465 0.1495 0.2501 1.5501 1.0050 1.0536 0.8189 —— ——
-Average 0.1606 0.2526 0.1531 0.2461 0.1559 0.2505 0.1563 0.2497 0.1537 0.2494 0.1628 0.2581 1.4888 0.9935 1.7061 0.9273 —— ——
-Exchange 96 14 0.0162 0.0830 0.0338 0.1331 0.0169 0.0834 0.0158 0.0807 0.0156 0.0825 0.0156 0.0799 2.5775 1.3205 3.1994 1.4322 1.1299 0.7540
-192 14 0.0180 0.0892 0.1413 0.2728 0.0182 0.0885 0.0174 0.0849 0.0171 0.0883 0.0166 0.0838 3.1525 1.4267 7.1367 2.9513 1.1609 0.7874
-720 14 0.0204 0.0992 0.0370 0.1441 0.0206 0.0999 0.0205 0.0964 0.0521 0.1683 0.0236 0.1092 2.2854 1.2949 1.0775 0.8829 1.4166 0.9583
-1440 14 0.0269 0.1202 0.0753 0.2084 0.1166 0.1882 0.0300 0.1203 0.0324 0.1320 0.1091 0.2550 2.8615 1.3872 7.9106 2.7104 1.5019 0.9877
-Average 0.0204 0.0979 0.0719 0.1896 0.0430 0.1150 0.0209 0.0955 0.0293 0.1178 0.0412 0.1320 2.7192 1.3573 4.8311 1.9942 1.3023 0.8719
-Weather 96 672 0.3529 0.3470 0.3691 0.3626 0.3432 0.3420 0.3383 0.3384 0.3553 0.3490 0.3585 0.3988 —— —— —— —— 0.4796 0.4392
-192 672 0.3401 0.3403 0.3602 0.3599 0.3285 0.3451 0.3245 0.3335 0.3398 0.3409 0.3502 0.3962 —— —— —— —— 0.5108 0.4526
-720 672 0.3142 0.3314 0.3630 0.3732 0.3337 0.3426 0.3200 0.3404 0.3179 0.3383 0.3266 0.3659 —— —— —— —— 0.4697 0.4512
-1440 672 0.3097 0.3304 0.3780 0.3877 0.3126 0.3349 0.3437 0.3596 0.3067 0.3318 0.3237 0.3595 —— —— —— —— 0.4765 0.4561
-Average 0.3292 0.3373 0.3676 0.3709 0.3295 0.3411 0.3316 0.3430 0.3299 0.3400 0.3398 0.3801 —— —— —— —— 0.4842 0.4498
-Wind 96 192 1.1953 0.7605 1.0926 0.7193 1.1628 0.7446 1.1471 0.7426 1.1796 0.7621 1.0439 0.7215 2.1130 1.0870 3.7312 1.3789 1.3909 0.8543
-192 192 1.1693 0.7601 1.1040 0.7268 1.1904 0.7582 1.1597 0.7477 1.1637 0.7639 1.0481 0.7211 1.8822 1.0471 3.7236 1.4758 1.4510 0.8885
-720 192 1.0969 0.7605 1.0309 0.7213 1.1576 0.7643 1.2393 0.7897 1.0970 0.7569 1.1007 0.7498 1.9993 1.1442 3.8525 1.4478 1.5825 0.9126
-1440 192 1.0434 0.7326 1.0339 0.7234 1.0880 0.7658 1.4168 0.8580 1.1244 0.7597 1.1375 0.7587 2.2880 1.1957 3.4273 1.3756 1.5407 0.9212
-Average 1.1262 0.7534 1.0653 0.7227 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
-Appendix B Training and Inference Algorithm
-Algorithms 1 and 2 illustrate the training algorithm and inference algorithm.
-$X_{L,s}^{0}=X_{L}^{0}-X_{L,t}^{0},$ (19)
-where the seasonal component $X_{L,s}^{0}$ is the residual after subtracting the trend.
-The denoising process starts by embedding $X_{T,s}^{k}$:
-$\mathbf{h}=\text{Emb}(X_{T,s}^{k}),$ (20)
-where $\text{Emb}(\cdot)$ is a convolutional MLP that maps the noisy sequence to a hidden dimension while capturing local temporal features.
-Sinusoidal positional encodings for diffusion step $t$ are integrated via Adaptive Layer Normalization (AdaLN):
-$\text{AdaLN}(\mathbf{h})=\alpha_{t}\cdot\text{LayerNorm}(\mathbf{h})+\beta_{t}.$ (21)
-where $\alpha_{t}$ and $\beta_{t}$ are learnable parameters projected from the diffusion step encoding.
-The predicted statistics are used in the denoising process via Gaussian sampling:
-$\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{T,i}+\hat{\sigma}_{T,i}\cdot z,$ (22)
-where $z\sim\mathcal{N}(0,I)$. This enables the model to incorporate stochastic variability while aligning with historical seasonal patterns.
-The denoised output $\hat{x}_{T,s}^{0}$ is combined with the conditional information through a weighted summation:
-$X_{\theta}\left(X_{T,s}^{k},t\mid\nabla_{\mathbf{x}_{T,s}}^{k-1}\right)=\rho_{1}\hat{x}_{T,s}^{0}+\rho_{2}\nabla_{\mathbf{x}_{T,s}}^{k-1}.$ (23)
-where $\rho_{1}$ and $\rho_{2}$ are learnable parameters that balance the denoised series and conditional information, ensuring effective seasonal modeling while maintaining statistical properties.
-The outputs from both pathways are combined via weighted summation to capture a wider range of trends:
-$f_{\theta}(x_{T,t}^{0})=\lambda_{1}T_{\text{origin}}(x_{T,t}^{0})+\lambda_{2}T_{\text{sqrt}}(x_{T,t}^{0}),$ (24)
-where $\lambda_{1}$ and $\lambda_{2}$ are learnable parameters that balance the contributions of the linear and root-based components.
-The final prediction $\hat{x}_{T}^{0}$ calculated by integrating the trend and seasonal components as follows:
-$\hat{x}_{T}^{0}=X_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{s,t}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\mid\nabla_{\mathbf{x}_{T,s}}^{k}\right)+f_{\theta}(x_{T,t}^{0}),$ (25)
-where $X_{\theta}$ represents CDSM and $f_{\theta}$ corresponds to PTM.
-The loss function in CDPM is derived from the ELBO:
-$L_{\text{ELBO}}=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\hat{x}_{T}^{0}\right\|^{2},$ (26)
-where $x_{T}$ and $\hat{x}_{T}$ represent the true value and the predicted value of the prediction window, respectively.
-Table 8: Performance comparison across different variants on ETTh2 and ETTm2.
-Dataset L T CDPM w/o CR PLR w/o INA w/o CR+PLR w/o DECOUPLING REV
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh2 96 192 0.3541 0.3837 0.3690 0.3889 0.3518 0.3819 0.3650 0.3873 0.3764 0.3940 0.3609 0.3874 0.3712 0.3947
-192 192 0.3476 0.3869 0.3480 0.3801 0.3466 0.3822 0.3403 0.3811 0.3426 0.3745 0.4039 0.4174 0.3878 0.4069
-720 192 0.3324 0.3811 0.3287 0.3732 0.3395 0.3806 0.3314 0.3756 0.3444 0.3782 0.4015 0.4258 0.3904 0.4153
-1440 192 0.3231 0.3787 0.3158 0.3701 0.3354 0.3856 0.3254 0.3819 0.3529 0.3862 0.3691 0.4118 0.3801 0.4139
-Average 0.3393 0.3826 0.3404 0.3781 0.3433 0.3826 0.3405 0.3815 0.3541 0.3832 0.3839 0.4106 0.3824 0.4077
-ETTm2 96 192 0.2445 0.3043 0.2461 0.3042 0.2453 0.3049 0.2465 0.3043 0.2585 0.3370 0.2435 0.3039 0.2456 0.3057
-192 192 0.2296 0.2959 0.2307 0.2953 0.2307 0.2970 0.2297 0.2961 0.2642 0.3417 0.2396 0.3013 0.2502 0.3102
-720 192 0.2200 0.2984 0.2279 0.3036 0.2379 0.3108 0.2192 0.2986 0.2331 0.3146 0.2488 0.3159 0.2517 0.3246
-1440 192 0.2269 0.3048 0.2232 0.3061 0.2659 0.3356 0.2268 0.3049 0.2295 0.3172 0.2632 0.3383 0.2737 0.3418
-Average 0.2303 0.3009 0.2320 0.3023 0.2450 0.3121 0.2305 0.3010 0.2463 0.3276 0.2488 0.3148 0.2553 0.3206
-Appendix C DERIVED ELBO for CDPM
-We begin by considering the standard diffusion model, which involves a forward process where data is gradually corrupted by noise, and a backward process to denoise it.
-In the forward diffusion process, the input $x^{0}$ is transformed into a noisy vector $x^{k}$ through the following steps:
-$q(x^{k}|x^{k-1})$ $=\mathcal{N}\left(x^{k};\sqrt{1-\beta_{k}}\,x^{k-1},\beta_{k}\,I\right),$ (27)
-where $\beta_{k}$ is the variance schedule, and $I$ is the identity matrix.
-By composing these steps, we can express the distribution of $x^{k}$ given $x^{0}$ as:
-$q(x^{k}|x^{0})$ $=\mathcal{N}\left(x^{k};\sqrt{\bar{\alpha}_{k}}\,x^{0},\left(1-\bar{\alpha}_{k}\right)I\right),$ (28)
-where $\bar{\alpha}_{k}=\prod_{i=1}^{k}(1-\beta_{i})$.
-Thus, $x^{k}$ can be directly expressed as:
-$x^{k}$ $=\sqrt{\bar{\alpha}_{k}}\,x^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (29)
-The backward denoising process is defined by a neural network that predicts the mean and variance:
-$p_{\theta}(x^{k-1}|x^{k})$ $=\mathcal{N}\left(x^{k-1};\mu_{\theta}(x^{k},k),\Sigma_{\theta}(x^{k},k)\right),$ (30)
-where $\Sigma_{\theta}(x^{k},k)$ is fixed as $\sigma_{k}^{2}I$, and $\mu_{\theta}(x^{k},k)$ is modeled by a neural network.
-The training objective minimizes the KL divergence between the forward and reverse processes:
-$L_{k}$ $=D_{\text{KL}}\left(q(x^{k-1}|x^{k},x^{0})\,\|\,p_{\theta}(x^{k-1}|x^{k})\right).$ (31)
-For stable training, $q(x^{k-1}|x^{k},x^{0})$ is simplified as:
-$q(x^{k-1}|x^{k},x^{0})$ $=\mathcal{N}\left(x^{k-1};\mu_{k}(x^{0},x^{k}),\tilde{\beta}_{k}I\right),$ (32)
-where $\mu_{k}(x^{0},x^{k})$ is expressed as:
-$\mu_{k}(x^{0},x^{k})$ $=\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\,x^{0}+\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\,x^{k}.$ (33)
-The training objective simplifies to:
-$L_{k}$ $=\frac{1}{2\sigma_{k}^{2}}\,\mathbb{E}_{k}\left[\left\|\mu_{k}(x^{0},x^{k})-\mu_{\theta}(x^{k},k)\right\|^{2}\right].$ (34)
-The mean of $x^{k-1}$, denoted $\mu_{x}(x_{\theta})$, is calculated as:
-$\mu_{x}(x_{\theta})$ $=\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\,x^{k}+\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\,x_{\theta}(x^{k},k).$ (35)
-The corresponding simplified loss is:
-$L_{x}$ $=\omega\mathbb{E}_{k,x^{0},\epsilon}\left[\left\|x^{0}-x_{\theta}(x^{k},k)\right\|^{2}\right],$ (36)
-where the parameter $\omega$ is a constant, and during training, it is typically set to 1.
-Now , we introduce normalize the $x_{L}$ and $x_{T}$, and then decompose the condition window $x_{L}^{0}$ and predicted series $x_{T}^{0}$ into its trend and seasonal components:
-$x_{L}^{0}$ $=x_{L,t}^{0}+x_{L,s}^{0}.$ (37)
-$x_{T}^{0}$ $=x_{T,t}^{0}+x_{T,s}^{0}.$ (38)
-We predict the trend component $x_{t}^{0}$ using $f_{\theta}$:
-$\hat{x}_{T,t}^{0}$ $=f_{\theta}(x_{L,t}^{0}).$ (39)
-Next, the seasonal component $x_{s}^{0}$ undergoes the diffusion process. The noisy version of the seasonal component, denoted by $r_{s}^{k}$, is obtained by adding noise to $x_{s}^{0}$:
-$x_{T,s}^{k}$ $=\sqrt{\bar{\alpha}_{k}}x_{T,s}^{0}+\sqrt{1-\bar{\alpha}_{k}}\,\epsilon.$ (40)
-CDPM iTransformer PatchMixer TSMixer PatchTST DLinear
-Figure 6: Comparison of CDPM and deterministic models on ETTh1 Dataset.
-CDPM iTransformer PatchMixer TSMixer PatchTST DLinear
-Figure 7: Comparison of CDPM and deterministic models on ETTh2 Dataset.
-We improve the denoising process by integrating local statistical features derived from the historical sequence $x_{L,s}^{0}$.
-The sequence is segmented into patches $P$, each with a length $H$, to capture complex seasonal variations.
-For each patch $P_{i}$, we compute the mean $\mu_{i}$ and variance $\sigma_{i}^{2}$, which encapsulate the local statistical properties:
-$\mu_{i}=\frac{1}{H}\sum_{h=1}^{H}x_{L,s,i,h}.$ (41)
-$\quad\sigma_{i}^{2}=\frac{1}{H}\sum_{h=1}^{H}\left(x_{L,s,i,h}-\mu_{i}\right)^{2}.$ (42)
-These computed statistics are utilized by a multi-layer perceptron (MLP) to predict the mean $\hat{\mu}_{i+1}$ and variance $\hat{\sigma}_{i+1}^{2}$ for the next patch:
-$\hat{\mu}_{i+1}=\text{MLP}(\mu_{i}).$ (43)
-$\quad\hat{\sigma}_{i+1}^{2}=\text{MLP}(\sigma_{i}^{2}).$ (44)
-The conditional input for denoising is then sampled as:
-$\nabla_{\mathbf{x}_{T,s}}^{k-1}=\hat{\mu}_{i+1}+\hat{\sigma}_{i+1}\cdot z,\quad z\sim\mathcal{N}(0,1).$ (45)
-This allows the seasonal component $\hat{x}_{T,s}^{0}$ to be predicted as:
-$\hat{x}_{T,s}^{0}=X_{\theta}(x_{T,s}^{k},k\mid\nabla_{\mathbf{x}_{T,s}}^{k}).$ (46)
-The final reconstruction of the time series is obtained by combining the predicted trend and seasonal components:
-$\hat{x}_{T}^{0}$ $=\hat{x}_{T,t}^{0}+\hat{x}_{T,s}^{0}.$ (47)
-To train the model, we minimize the mean squared error (MSE) between the true and predicted values for both components. The individual losses are defined as follows:
-The trend loss measures the error between the true and predicted trend components:
-$L_{\text{trend}}$ $=\mathbb{E}_{x_{T}^{0}}\left[\left\|x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right\|^{2}\right].$ (48)
-The seasonal loss captures the error between the true and predicted seasonal components:
-$L_{\text{seasonal}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left[\left\|x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right\|^{2}\right].$ (49)
-We introduce a cross-term loss to account for the interaction between the trend and seasonal components.
-This term ensures the model considers any dependency between these components during prediction.
-The cross-term loss is defined as:
-$L_{\text{cross}}$ $=2\mathbb{E}_{x_{T}^{0}}\left[\left(x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right)\cdot\mathbb{E}_{\epsilon}\left[\left(x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right)\mid x_{T}^{0}\right]\right]$ (50)
-$=2\mathbb{E}_{x_{T}^{0},\epsilon}\left[\left(x_{T,t}^{0}-\hat{x}_{T,t}^{0}\right)\cdot\left(x_{T,s}^{0}-\hat{x}_{T,s}^{0}\right)\right].$ (51)
-The total loss function is the combination of the trend, seasonal, and cross-term losses:
-$L_{\text{mse}}$ $=L_{\text{trend}}+L_{\text{seasonal}}+L_{\text{cross}}.$ (52)
-Expanding the total loss, we obtain:
-$L_{\text{mse}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}^{0}-\left(\hat{x}_{T,t}^{0}+\hat{x}_{T,s}^{0}\right)\right\|^{2}.$ (53)
-Finally, we denormalize and integrate the diffusion model into the seasonal component, the loss function becomes:
-$L_{\text{mse}}$ $=\mathbb{E}_{x_{T}^{0},\epsilon}\left\|x_{T}-\hat{x}_{T}\right\|^{2}.$ (54)
-This total loss function balances the contributions of the trend and seasonal components while accounting for their interactions.
-By minimizing this loss, the model learns to reconstruct the time series with greater accuracy.
-CDPM CSDI Diffusion-TS
-Figure 8: Comparison of CDPM and DDPM Models on ETTh1 Dataset.
-CDPM CSDI Diffusion-TS
-Figure 9: Comparison of CDPM and DDPM Models on ETTh2 Dataset.
-Appendix D EXPERIMENTAL RESULTS
-D.1 Overall Performance Results
-The experimental results presented in Table 7 provide compelling evidence of the effectiveness of the Conditional Denoising Polynomial Modeling (CDPM) framework for time series forecasting.
-This advantage stems from CDPM’s decoupling strategy, which separates trend and seasonal components for more precise modeling of temporal patterns.
-This validates the framework’s theoretical contributions, where decoupling was hypothesized to improve accuracy and interpretability, now confirmed by the experimental results.
-D.2 Overall Ablation Experiments
-We conduct a comprehensive ablation study to evaluate the impact of key components in the proposed CDPM architecture.
-Specifically, we examine the role of the condition information injection module, the PTM, and instance normalization in handling the complex temporal dynamics present in time series data.
-The study is performed on the ETTh2 and ETTm2 datasets, and the results are presented in Table 8.
-D.3 Case Study
-As shown in Fig.6 and Fig.7, we compare CDPM with deterministic and diffusion-based models on the ETTh1 and ETTh2 datasets.
-With a history length of 192, CDPM consistently outperforms other models in capturing trends and seasonal variations.
-On ETTh1, CDPM closely aligns with the ground truth, excelling in modeling complex seasonal fluctuations and trend shifts.
-In contrast, deterministic models like PatchMixer and TSMixer fail to capture rapid trend changes and local seasonal variations, while Autoformer and DLinear produce overly smooth or delayed predictions.
-CDPM strikes an optimal balance in handling complex time series dynamics.
-In Fig.8 and Fig.9, CDPM is compared with diffusion-based models like CSDI and Diffusion-TS.
-CDPM offers smoother, more accurate forecasts, while Diffusion-TS shows higher variability and lower reliability, especially on ETTh2.
-By avoiding over-smoothing, CDPM better handles complex temporal structures.
-These results underscore CDPM’s ability to capture both trends and fluctuations, outperforming deterministic models and delivering more reliable forecasts than diffusion-based models.
-The advantages of CDPM’s decoupling strategy and specialized modules are clear.
+Extensive experiments on real-world datasets validated the effectiveness of the proposed CDPM.
+Acknowledgements
+This research was supported by grants from the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), the National Natural Science Foundation of China (62337001), the Key Technologies R&D Program of Anhui Province (No.
+202423k09020039), USTC Research Funds of the Double First-Class Initiative (No.
+YD2150002501), and the Fundamental Research Funds for the Central Universities (No.
+WK2150110032).

### v7

@@
 Extensive experiments conducted on six benchmarks demonstrate the effectiveness of our framework, highlighting the potential of combining probabilistic and deterministic models.
-[^2] Our code is available at https://github.com/zjt-gpu/CDPM.
+Our code is available at https://github.com/zjt-gpu/CDPM.
 1 Introduction
@@
 Diffusion-based methods Lee et al.
-(2024) have shown significant potential in time series forecasting.
-We argue that these approaches are critically limited by the naive application of noise addition and denoising processes, which fails to account for the distinct dynamics of different patterns.
-Specifically, noise is uniformly added across the entire time series, and denoising is then applied indiscriminately, treating all temporal components equally.
+(2024) have shown great promise in time series forecasting.
+However, we argue that their naive application of noise addition and denoising overlooks the heterogeneous nature of temporal dynamics.
+In particular, these methods uniformly add noise across the entire sequence and apply denoising without distinguishing between different temporal components.
 However, time series inherently exhibits diverse dynamics, including trends and seasonal patterns, which require different treatment.
 Trends patterns, reflecting gradual shifts over extended periods, and seasonal patterns, characterized by periodic fluctuations, possess different structural properties.
-This strategy of uniform noise addition and denoising fails to account for these differences, potentially distorting key components, particularly intrinsic trends and fluctuations, which are crucial for accurate forecasting.
+These distinct patterns require differentiated treatment.
+Such indiscriminate processing can distort essential features, especially intrinsic trends and fluctuations, thereby degrading forecasting performance.
 As shown in Fig.1, this indiscriminate approach disrupts critical trend information during denoising, hindering trend recovery and reducing accuracy.
 Typically, a single model often struggles to capture the subtle and dynamic interplay between trends and variations, which can significantly affect the model’s overall forecasting ability.
-Diffusion models, which excel at modeling complex dynamic processes, capturing high-frequency fluctuations, and quantifying uncertainty, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series.
+Diffusion models, which excel at modeling complex dynamic processes, are more suitable for handling components characterized by large fluctuations and high-frequency variations in time series.
 In contrast, deterministic models, such as models based on linear layers, learn features from historical patterns and are particularly effective at identifying regular patterns, making them better suited for modeling components with stable trends.
@@
 where the seasonal component $X_{L,s}^{0}$ is the residual after subtracting the trend. Thus, we obtain the trend component $X_{L,t}^{0}$ and seasonal component $X_{L,s}^{0}$ of historical window.
-Similarly, we compute $X_{T,t}^{0}$ and $X_{T,s}^{0}$ for predicted window. Decomposition operation isolates the distinct characteristics of time series.
+Similarly, we compute $X_{T,t}^{0}$ and $X_{T,s}^{0}$. Decomposition operation isolates the distinct characteristics of time series.
 4.3 Conditional Denoising Seasonal Module
-Modeling fluctuating seasonal components in time series is challenging due to their non-stationarity and intricate patterns.
-We propose CDSM, based on a denoising diffusion model, a popular paradigm in probabilistic modeling.
+Modeling fluctuating seasonal components in time series is challenging due to their intricate patterns.
+We propose CDSM, based on a denoising diffusion model.
 By incorporating historical statistical properties as conditional inputs, CDSM captures seasonal patterns more effectively.
@@
 $\sigma_{L,i}^{2}=\frac{1}{H}\sum_{h=1}^{H}\left(x_{L,s,i,h}-\mu_{L,i}\right)^{2},$ (7)
-where, for each patch $P_{L,i}$, $\mu_{L,i}$ represents the local mean and $\sigma_{L,i}^{2}$ denotes the variance.
-The historical data is divided into patches $P$ of length $H$, with each patch capturing localized seasonal signals.
+where, for each patch $P_{L,i}$, $\mu_{L,i}$ represents the local mean and $\sigma_{L,i}^{2}$ denotes the variance. The historical data is divided into patches $P$ of length $H$.
 These statistics capture seasonal variability. The predicted mean $\hat{\mu}_{T,i}$ and variance $\hat{\sigma}_{T,i}^{2}$ for the target window $P_{T,i}$ are computed using two MLPs:
@@
 4.4 Polynomial Trend Module
-Different trends exhibit distinct patterns of progression, including periods of growth and decline, making it essential to capture a broader range of trend behaviors.
+Different trends follow distinct trajectories, including phases of growth and decline, highlighting the need to capture diverse trend behaviors.
 By leveraging Cover’s theorem Cover (1965), we project smooth time series patterns into a polynomial space, facilitating the extraction of both linear and nonlinear trends.
@@
 (2021) is employed to better capture these diverse trend variations.
-To fully harness polynomial modeling, we introduce a square root transformation as a nonlinear transformation within the polynomial framework.
-The square root transformation effectively smooths extreme values, enhancing the framework’s flexibility in capturing a wide range of trend patterns.
+To fully harness polynomial modeling, we introduce a square root transformation to smooth series and enhance the framework’s flexibility in capturing diverse trend patterns.
 As a deterministic model, PTM consists of two complementary pathways. The first models stable trends with a linear layer applied directly to the historical trend series:
@@
 Component Reconstruction.
-During the training and inference phases, after the trend component $\hat{x}_{T,t}^{0}$ is predicted by the PTM and the seasonal component $\hat{x}_{T,s}^{0}$ is recovered by the CDSM.
-The final prediction $\hat{x}_{T}^{0}$ is calculated by integrating the trend and seasonal components as follows:
+The final prediction $\hat{x}_{T}^{0}$ is calculated by integrating the trend $\hat{x}_{T,t}^{0}$ and seasonal components $\hat{x}_{T,s}^{0}$ as follows:
 $\hat{x}_{T}^{0}=X_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{s,t}^{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\mid\nabla_{\mathbf{x}_{T,s}}^{k}\right)+f_{\theta}(x_{T,t}^{0}),$ (15)
@@
 This formulation combines the strengths of both probabilistic diffusion models and deterministic linear models, leading to more accurate forecasts.
-5 Experiments
-5.1 Experimental Settings
-Table 1: Overview of the dataset characteristics. Here, $T$ denotes the target series used for prediction.
 Dataset ETTh1 ETTm1 Wind Exchange Weather Electricity
@@
 T (Steps) 168 192 192 14 672 168
-Table 2: The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined.
-CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
+Table 1: Overview of the dataset characteristics.
+5 Experiments
+5.1 Experimental Settings
 Dataset T CDPM iTransformer Patchmixer TSMixer PatchTST DLinear Diffusion-TS D3VAE CSDI
@@
 Wind 192 1.1262 0.7534 1.0653 0.7227 1.1497 0.7582 1.2407 0.7845 1.1411 0.7607 1.0826 0.7378 2.0706 1.1185 3.6837 1.4195 1.4913 0.8942
+Table 2: The forecasting results on eight datasets, with the best results highlighted in bold and the second-best underlined.
+CSDI encounters out-of-memory issues on the Electricity and Diffusion-TS datasets, while D3VAE runs out of memory on the Weather dataset.
 Datasets.
@@
 CDPM’s strength lies in its Polynomial Trend Module (PTM) and Conditional Denoising Seasonal Module (CDSM), which capture complex trends and seasonal dynamics.
-On high-dimensional datasets like Electricity and Weather, CDPM maintains superior accuracy.
 Unlike CSDI and Diffusion-TS, which struggle with long forecast horizons, CDPM handles long-term dependencies effectively.
@@
 Overall, CDPM shows strong generalization across various datasets.
-In summary, CDPM’s decoupling strategy enhances performance, confirming the theoretical hypothesis that it improves accuracy and interpretability over traditional and diffusion-based models.
 Ablation Study.
-Table 3: Ablation study results on ETTh2 and ETTm2.
 Method ETTh2 ETTm2
@@
 CDPM4 0.3442 0.3807 0.2335 0.3015
+Table 3: Ablation study results on ETTh2 and ETTm2.
 The ablation experiments highlight the importance of each component in the CDPM framework, especially for handling complex temporal dynamics in datasets with varying granularities like ETTh2 and ETTm2.
@@
 Removing the condition information injection module reduces performance, with a larger impact on ETTm2, where external factors influence intricate temporal patterns.
-Replacing the PTM with a linear layer causes a significant performance drop, particularly on ETTm2, underscoring the need for non-linear trend modeling.
-PTM is crucial for capturing long-term trends and periodic behaviors.
-Eliminating instance normalization leads to the largest performance decline, especially on ETTm2, highlighting its role in stabilizing training for high-resolution datasets.
-On ETTh2, instance normalization slightly improves performance, indicating its lesser importance when trends and seasonality are clearer.
+Replacing the PTM with a linear layer causes a significant performance drop, which is crucial for capturing long-term trends and periodic behaviors.
+Eliminating instance normalization leads to the largest performance decline, highlighting its role in stabilizing training for high-resolution datasets.
 Removing both the condition information injection module and PTM further degrades performance, emphasizing the necessity of both components.
-Figure 3: Parameter sensitivity analysis of CDPM: hidden dimension (left), patch length (middle), and predicted length (right).
-Figure 4: (a) Visual comparison of CDSM (left) and PTM (right) in modeling seasonal components. (b) Visual comparison of CDSM (left) and PTM (right) in modeling trend components.
+Method ETTh2 ETTm2
+MSE MAE MSE MAE
+CDPM 0.3393 0.3826 0.2303 0.3009
+CDPM5 0.3839 0.4106 0.2518 0.3157
+CDPM6 0.3541 0.3832 0.2452 0.3122
+Table 4: Decoupling validation study results on ETTh2 and ETTm2.
 Evaluation of the Decoupling Strategy.
@@
   • CDPM6: In this variant, PTM and CDSM are reversed, with PTM predicting the seasonal component and CDSM predicting the trend component.
-The decoupling strategy, which separates trend and seasonal patterns, significantly enhances performance, with our framework CDPM outperforming all ablated models.
-The improvement is most pronounced on ETTh2, where clear trend-seasonal separation aids effective modeling.
-On ETTm2, the gain is more modest due to finer granularity and overlapping patterns, but CDPM still outperforms the ablated models, demonstrating its adaptability.
+The decoupling strategy that separates trend and seasonal patterns markedly improves performance.
 Applying our framework directly underperforms due to its inability to separate trend and seasonal components, while reversing PTM and CDSM shows slight improvements but still falls short of our framework, highlighting the importance of proper module assignment.
-These results emphasize the need to tailor modules to capture distinct trend and seasonal patterns, with decoupling especially beneficial for datasets with clear temporal structures.
-For finer-grained datasets like ETTm2, the gains from decoupling are more modest, but our framework consistently outperforms the ablated models, showcasing its robustness across datasets with varying granularities.
-Table 4: Decoupling validation study results on ETTh2 and ETTm2.
-Method ETTh2 ETTm2
-MSE MAE MSE MAE
-CDPM 0.3393 0.3826 0.2303 0.3009
-CDPM5 0.3839 0.4106 0.2518 0.3157
-CDPM6 0.3541 0.3832 0.2452 0.3122
+These results emphasize the need to tailor modules for distinct patterns, with decoupling especially beneficial for datasets with clear temporal structure.
+Figure 3: Parameter sensitivity analysis of CDPM: hidden dimension (left), patch length (middle), and predicted length (right).
+Figure 4: (a) Visual comparison of CDSM (left) and PTM (right) in modeling seasonal components. (b) Visual comparison of CDSM (left) and PTM (right) in modeling trend components.
 PTM and CDSM Modeling Analysis.
-To compare the performance of the specially designed components, PTM and CDSM, in modeling the trend and seasonal components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168.
-The results highlight the effectiveness of tailored modeling strategies for each component and the benefits of decoupled modeling in improving forecasting accuracy.
+To evaluate the effectiveness of PTM and CDSM in modeling different components of the ETTh1 dataset, we conducted experiments with an input length of 192 and a prediction horizon of 168.
+The results highlight the effectiveness of tailored modeling strategies for each component.
 As shown in Figure 4 (a), PTM excels in capturing intrinsic and gradual trends due to its linear structure, accurately modeling patterns and tracking the overall direction of the time series.
@@
 For the seasonal component, which exhibits complex fluctuations, the seasonal module demonstrates superior performance, as shown in Figure 4 (b).
-CDSM effectively captures sharp seasonal changes, while PTM focuses on underlying regularities in the fluctuations, though it may sometimes miss rapid shifts.
 These results validate each module’s modeling capabilities for its designated component.
@@
 Similarly, $H$ significantly affects performance, with $H=8$ capturing both local dependencies and global patterns, whereas larger values introduce redundancy and reduce efficiency.
-Finally, our framework performs consistently across prediction horizons, with the trend module capturing long-term structure and the seasonal module adapting to fluctuations, ensuring reliable forecasts.
-Overall, the analysis confirms the balance between complexity and scalability in our framework, making it suitable for diverse forecasting tasks.
+Finally, our framework performs consistently across prediction horizons.
 Figure 5: Visualization of trend (left) and seasonall (right) components of joint modeling for the electricity dataset.